### v1

# A Super Version of a theorem of Fricke-Klein

Abstract.
In this paper, we study the character variety of the supergroup $\operatorname{OSp}(1|2)$ in the specific case of the free group on two letters as a first step towards an algebraic treatment for the character variety of supergroups. One can interpret this as an analogue of a result of Fricke and Klein, who studied this in the case of $\operatorname{SL}(2,\mathbb{C})$. We give a count of independent generators of the ring of invariants.
1. Introduction
A result by Fricke and Klein [8] and apparently Vogt [25] states, that the $SL(2,\mathbb{C})$-invariant functions on $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ are given by the ring
(1) $\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)].$
In the context of geometric invariant theory, we can interpret this result as
$\operatorname{Spec}(SL_{2})\times\operatorname{Spec}(SL_{2})$$\operatorname{Spec}(\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)])\cong\mathbb{C}^{3}$$\scriptstyle{\pi}$
i.e. a quotient variety. Similarly, another reinterpretation of this result , which is even more geometric, is that it is equivalent to the fact that the character variety of the free group on two letters and $SL(2,\mathbb{C})$ is isomorphic to $\mathbb{C}^{3}$, i.e.
(2) $\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}.$
In this paper we are aiming to generalize the third interpretation of this result, to the super world. The motivation of this result comes from the fact, that $\operatorname{OSp}(1|2)$ is the super generalization of $\operatorname{SL}(2,\mathbb{C})$ and takes a similar place in the theory of super Riemann surfaces, as $\operatorname{SL}(2,\mathbb{C})$ does in the classical theory. On the real analytic side of this story, we get something called super Teichmüller space, where each connected component of the space is homeomorphic to $\mathbb{R}^{6g-6|4g-4}/\mathbb{Z}_{2}$ [2] [19], where $\mathbb{Z}_{2}$ only acts on the odd coordinates.
By ”explicit” computation, we can show that there is a deviation from the classical case and we do not get a variety of the expected dimension $3|2$. A problem of supergeometry is that most of the time the super version of the proof works analogously without adjusting anything. We want to close the gap of the unknown by writing down the super versions of the proof, even if nothing changes, so the reader does not have to mentally go through the classical construction and reason, that it works accordingly in the super case.
Organization of the paper. In section 2 we will give a short exposition on the result of Fricke and Klein. In section 3 introduce basic notions of supergeometry, algebraic supergroups and invariant theory to understand our main result. In section 4 we introduce the character variety, which is the geometric interpretation of our result, where we can also easily see that this definition adapts to super. In section 5 we prove our main result
Theorem 1.1.
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ are generated by 7 traces.
Acknowledgments. I would like to thank my advisor Paul Norbury for his constant guidance and confidence in my work. I would also like to thank Miles Koumouris, Gustav Berth and Lior Yanovski for useful discussion.
I am supported by the University of Melbourne Graduate Research Scholarship.
2. Fricke-Klein
In this section, we give an elementary proof of the result of Fricke and Klein and apparently Vogt. In more recent times, Goldman reproved this theorem in [10], where he used the Cayley-Hamilton theorem as in an essential way in the proof. In this version, we forego the usage of the Cayley-Hamilton theorem and instead make use of a simultaneous conjugation of a pair of matrices into upper and lower triangular forms. Before we start proving the result of Fricke and Klein, let us first state the aforementioned triangulation lemma.
Lemma 2.1.
Let $A_{0}\in\operatorname{SL}(2,\mathbb{C})$ and $B_{0}\in\operatorname{SL}(2,\mathbb{C})$. Then one can always find an element $g\in\operatorname{SL}(2,\mathbb{C})$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
(3) $gA_{0}g^{-1}=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}\quad,\quad B_{0}=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}$
Proof.
This splits into two cases. $A$ diagonalizable or not. If $A$ is diagonalizable, then it must mean that for its eigenvalues $\mu,\mu^{-1}$ that $\mu\neq\mu^{-1}$. Consider
$C=\begin{pmatrix}0&-x^{-1}\\ x&y\\ \end{pmatrix}$
and compute its action by conjugation on A and B
$CAC^{-1}=\begin{pmatrix}\mu^{-1}&0\\ xy(\mu-\mu^{-1})&\mu\end{pmatrix}\quad,\quad CBC^{-1}=\begin{pmatrix}-x^{-1}yc+d&x^{-2}c\\ xya+y^{2}c-x^{2}b-xyd&a+x^{-1}yc\end{pmatrix}$
To set the bottom left corner of the first matrix to 1, we want $x=y^{-1}(\mu-\mu^{-1})^{-1}$. By substituting the $x$, the bottom left corner is equal to
$cy^{2}+\frac{a-d}{\mu-\mu^{-1}}-\frac{b}{\mu-\mu^{-1}}y^{-2},$
which we want to set to zero. So we can solve a polynomial of degree 4, to get an upper triangular matrix for $B$ and a lower triangular matrix $A$, by simultaneous conjugation.
Let $A$ not diagonalizable, then it must admit Jordan normal form
(4) $A=\begin{pmatrix}\mu&1\\ 0&\mu\end{pmatrix}$.
$A\in\operatorname{SL}(2,\mathbb{C})$, therefore $\mu=1$, which is a case we can treat by hands. Consider
$\begin{pmatrix}x&y\\ 0&z\end{pmatrix}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}z&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}1&x^{2}\\ 0&1\end{pmatrix}$
On the other hand for the arbitrary matrix we get
$\begin{pmatrix}x&y\\ 0&z\end{pmatrix}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\begin{pmatrix}z&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}z(xa+yc)&-y(xa+yc)+x(xb+yd)\\ z^{2}c&-yzc+xzd\end{pmatrix}$
Also notice, that
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}a&b\\ 0&d\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}=\begin{pmatrix}a&0\\ -b&d\end{pmatrix}$
First, we set $x=i$, then we choose a solution for the equation y(xa+yc)+x(xb+yd)-y(xa+yc)+x(xb+yd), which we can do over $\mathbb{C}$. Act by rotation by 90 degree matrix. and we get the pair
$A=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\quad,\quad B=\begin{pmatrix}z(xa+yc)&z^{2}c\\ 0&-yzc+xzd\end{pmatrix}$,
which is in the desired form. So, indeed we can always parametrize into the simultaneous triangular form. ∎
Remark 2.2.
There exists an analogous result for $\operatorname{OSp}(1|2)$, which will be proven later.
We are now ready to give a simple proof of the theorem mentioned in the introduction.
Theorem 2.3.
Let $A,B\in\operatorname{SL}(2,\mathbb{C})$ then all conjugation-invariant polynomial functions $f:SL(2,\mathbb{C})\times SL(2,\mathbb{C})\rightarrow\mathbb{C}$ are generated by $\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)$.
Proof.
Consider the coordinate ring of $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ denoted by $\mathbb{C}[G\times G]$ and its subring of $G$-invariant polynomial functions, denoted by $\mathbb{C}[H]$. We want to prove
[…]
where $X=\operatorname{tr}(A)$, $Y=\operatorname{tr}(B)$ and $Z=\operatorname{tr}(AB)$. First of all note, that if $f\in\mathbb{C}[H]$, then $f$ is a polynomial function in the entries of the two matrices $(\xi,\eta)\in SL(2,\mathbb{C})\times SL(2,\mathbb{C})$. We will denote it $f(A,B)$. Furthermore, by the triangulation lemma we can bring them into the form
$A=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}\quad,\quad B=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}$
This parametrization means, that we can view $f\in\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$, which makes $\mathbb{C}[H]$ into a subring by the obvious inclusion map
$\mathbb{C}[H]\rightarrow\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa].$
We can also define the map
$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$ $\rightarrow\mathbb{C}[X,Y,Z,\Delta_{X},\Delta_{Y}]$
$\lambda$ $\mapsto\frac{X+\Delta_{X}}{2}$
$\lambda^{-1}$ $\mapsto\frac{X-\Delta_{X}}{2}$
$\mu$ $\mapsto\frac{Y+\Delta_{Y}}{2}$
$\mu^{-1}$ $\mapsto\frac{Y-\Delta_{Y}}{2}$
$\kappa$ $\mapsto Z-\frac{X+\Delta_{X}}{2}\frac{Y+\Delta_{Y}}{2}-\frac{X-\Delta_{X}}{2}\frac{Y-\Delta_{Y}}{2},$
where $\Delta_{X}=\sqrt{X^{2}-4}$ and $\Delta_{Y}=\sqrt{Y^{2}-4}$. We can easily see that
[…]
If we now consider a polynomial $f\in\mathbb{C}[X,Y,Z]\subset\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$, it is easy to see it is $G$-invariant, thus $f\in\mathbb{C}[H]$. Notice that
$\lambda-\lambda^{-1}$ $\mapsto\Delta_{X}$
$\mu-\mu^{-1}$ $\mapsto\Delta_{Y}.$
So it turns out, that the statement is even stronger, we have
[…]
Now we define a map $\psi$, by the composition of the solid arrows in the following diagram
$\mathbb{C}[H]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\mathbb{C}[X,Y,Z]$$\mathbb{C}[X,Y,Z,\Delta_{X},\Delta_{Y}]$$\scriptstyle{\eqqcolon\psi}$
which is surjective as argued earlier. Now to prove it is injective, we again make use of our identifications
$\mathbb{C}[H]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\mathbb{C}[X,Y,Z]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\scriptstyle{\operatorname{id}}$
Now consider the image of a $G$-invariant polynomial $g$. We have $\psi(g(\xi,\eta))=f(X,Y,Z)=0$. This is equivalent to
$\sum_{ijk}\alpha_{ijk}X^{i}Y^{j}Z^{k}=\sum_{ijk}\alpha_{ijk}(\lambda+\lambda^{-1})^{i}(\mu+\mu^{-1})^{j}(\lambda\mu+\lambda^{-1}\mu^{-1}+\kappa)^{k}=0,$
which we expand in the form
$\sum_{ijk}\alpha_{ijk}\sum_{l}^{i}\sum_{m}^{j}\sum_{k_{1}+k_{2}=k_{3}=n}^{k}\binom{i}{l}\binom{j}{m}\binom{k}{k_{1},k_{2},k_{3}}\lambda^{i-2l+k_{1}-k_{2}}\mu^{j-2m+k_{1}-k_{2}}\kappa^{k_{3}}$
On the other hand we have the $G$-invariant polynomial
$\sum_{ijkpq}\beta_{ijkpq}\lambda^{i-j}\mu^{k-p}\kappa^{q}=0,$
which after relabeling turns into
$\sum_{rst}\beta_{rst}\lambda^{r}\mu^{s}\kappa^{t}=0.$
We can clearly see, every $\beta_{rst}$ is just a sum of the $\alpha_{ijk}$ By our condition we know all coefficients $\alpha_{ijk}$ must vanish, since there is no relation between $X,Y,Z$. Therefore $\beta_{rst}=0$. Thus $g=0$ and injectivity follows. In total we now have
$\mathbb{C}[G\times G]^{G}=\mathbb{C}[H]\cong\mathbb{C}[X,Y,Z]$
and on the level of spaces we have
$\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}$
3. Preliminaries
3.1. Supergeometry
Here we want to introduce basic notions of supergeometry following [11] and [3]. Generally, we will simply endow everything with a $\mathbb{Z}_{2}$-grading.
Definition 3.1.
Let $V=V_{0}\oplus V_{1}$ be a $\mathbb{Z}_{2}$-graded vector space, then it is a super vector space
Definition 3.2.
A superscheme $X=(|X|,\mathcal{O}_{X})$ is a locally ringed space, where $|X|$ is a topological space and $\mathcal{O}_{X}$ is a sheaf of $\mathbb{Z}_{2}$-graded commutative rings.
Remark 3.3.
We can interpret the odd coordinates geometrically as fat points. To elaborate a little bit on that:
Odd coordinates vanish if you square them, so if we think of a scheme over $\mathbb{C}$ of odd dimension $0|1$, we can take a look at the coordinate ring $\mathbb{C}[\theta]$ which is roughly the same as $\mathbb{C}[\epsilon]/\epsilon^{2}$, where $\theta$ is an odd coordinate, while $\epsilon$ is an even one.
By this analogy, we can think of odd coordinates geometrically as smearing out the point or a first order taylor expansion around the origin.
Example 3.4.
$\mathbb{C}^{n|m}$.
Example 3.5.
$\mathbb{P}^{n|m}$. Here we want to construct the super version of projective space. As underlying topological space we use $\mathbb{P}^{n}$, which is a collection of $U_{i}=\{[z_{0}:...:z_{n}]|z_{i}\neq 0\}$. These subsets are isomorphic to $\mathbb{C}^{n+1}$, because $z_{i}$ is always non zero, so we can divide by $z_{i}$ and get $U_{i}=\{[\frac{z_{0}}{z_{i}}:,...,1,...,\frac{z_{n}}{z_{i}}]\}$ . So if we take the corresponding locally free sheaf of rank m, then the associated supermanifold $(U_{i},\mathcal{O}_{U_{i}})$ is going to be $\mathbb{C}^{n|m}$. We can also define an isomorphism on the set $U_{i}\cap U_{j}$. We can choose n even generating sections of $U_{i}$, called $u_{i,a}$, and m odd generating sections called $\xi_{i,b}$. So we write:
(5) $\phi_{ij}(u_{i,0},...,\hat{u}_{i,i},...,u_{i,n},\xi_{i,1},...,\xi_{i,m})\mapsto(\frac{u_{i,0}}{u_{i,j}},..,\frac{1}{u_{i}},...,\frac{u_{i,n}}{u_{i,j}},\frac{\xi_{i,1}}{u_{i,j}},...,\frac{\xi_{i,m}}{u_{i,j}})$
We can also do this the other way around and can see, that the transition function on the intersection is $\frac{u_{i,i}}{u_{i,j}}$. One can also easily see, that these $\phi_{ij}$ satisfy the cocycle condition. So we can glue all these sections together and define a sheaf $\mathcal{O}_{\mathbb{P}^{n|m}}$. Together with the underlying topological space we have defined the super projective space $(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n|m}})$.
Definition 3.6.
The reduced space of a superscheme $X$ is defined as
(6) $X\coloneqq(|X|,\mathcal{O}_{X}/J_{X})=(|X|,\mathcal{O}_{red})$
where $J_{X}$ is the sheaf of ideals generated by the nilpotent sections of $\mathcal{O}_{X}$.
It is important to note, that $\mathcal{O}_{X}/J\coloneqq a(U\mapsto\mathcal{O}_{X}(U)/J(U))$ is the sheafification. However, we still get an isomorphism on stalks between the presheaf and its sheafification and hence we can use results known from commutative algebra with impunity. This notion is useful, because when we try to generalize our concepts to try to describe more abstract physics, we want to ensure that it does reduce to our known description of phenomena up to that point. However, there is generally an obstruction to being able to do this nicely. Ideally, we want a global section to exist, which is the projection onto the non-super space. So, we will introduce the notion of projected and split superscheme.
Definition 3.7.
A superscheme X is called projected if the structural sequence
[…]
splits.
As one might guess, this implies the existence of non-projected superscheme, but contrary to what our intuition might lead us to conclude a lot of supermanifolds are non-projected in even quite simple examples.
Remark 3.8.
In the category of supermanifolds, it turns out that every real supermanifold is split, due to Batchelors theorem (in a non-natural way). So if we want to look for non-projected objects, we must look at complex supermanifolds.
Definition 3.9.
Let X be a superscheme of odd dimension m, then there exists a $J_{X}$ adic filtration of $\mathcal{O}_{X}$
(7) $J_{X}^{0}\supset J_{X}^{1}\supset...\supset J_{X}^{m+1}=0.$
We can now define the sheaf
(8) $Gr(\mathcal{O}_{X})\coloneqq J_{X}^{0}/J_{X}^{1}\oplus J_{X}^{1}/J_{X}^{2}\oplus...\oplus J_{X}^{m-1}/J_{X}^{m}\oplus J^{m}_{X}$
The split superscheme associated to X is defined as
(9) $Gr(X)\coloneqq(|X|,Gr(\mathcal{O}_{X})).$
Definition 3.10.
A superscheme $X$ is called split, if there exists an isomorphism $X\cong Gr(X)$.
Lemma 3.11.
If a supermanifold X of odd dimension $m$ is split, it is projected.
Remark 3.12.
The terminology in supergeometry is not yet set in stone. Some authors refer to this property as split and some authors call this property being a graded supermanifold. We remind the reader to always carefully check, what is meant by the term graded or split when encountering supergeometry.
Remark 3.13.
A more categorical way and maybe more uniform way to view supergeometry in the context of ”prefix” algebraic geometry. Let $\operatorname{Ab}_{\mathbb{Z}_{2}}$ be the category of $\mathbb{Z}_{2}$-graded abelian groups, which is symmetric monoidal, then a superring $A$ is an commutative algebra/commutative monoid object inside of $\operatorname{Ab}_{\mathbb{Z}_{2}}$, which is usually denoted $\operatorname{CAlg}(\operatorname{Ab}_{\mathbb{Z}_{2}})\cong\operatorname{SRings}$.
3.2. Supergroups
Here we will introduce algebraic supergroups following [16] [15], the analogue of algebraic groups in supergeometry. Algebraic supergroups naturally appear, when dealing with supersymmetry related models in physics, however a more mathematics oriented answer as in how they appear is that representations of supergroups naturally appear in Delignes theorem on tensor categories [6]. One heuristic way of why they appear is that the category of vector spaces is not algebraically closed so to say, there is a natural $\mathbb{Z}_{2}$-action on the category of vector spaces and taking its closure with respect to that will result in the category of super vector spaces. So when you are trying to turn a tensor category into a category of representations you will see that if you restrict to ordinary vector spaces that you simply do not have enough points in usual setting one would move from $\mathbb{R}$ to $\mathbb{C}$, while in this case on moves from $\mathrm{Vect}_{k}$ to $\mathrm{sVect}_{k}$ for $k$ algebraically closed and of characteristic 0.
In standard literature schemes are described as locally ringed spaces equipped with the appropriate sheaf, that extracts the algebraic structure. This formalism allows us to easily enter the realm of supergeometry, by replacing our sheaves with $\mathbb{Z}_{2}$-graded sheafs. Similarly, algebraic groups can be studied as a locally ringed space. In the affine case, to make the group structure compatible/easily translatable through our functor Spec, we should describe it as a set of arrows that we can invert, when applying Spec. This procedure results in the following
Definition 3.14.
Let $G$ be superscheme and consider $m:G\times G\rightarrow G$ to be a morphism of schemes. Furthermore, if there exists morphisms of superschemes
$e:*\rightarrow G\quad,\quad\operatorname{inv}:G\rightarrow G$
such that the following diagrams commute
[…]
then we call $G$ a group superscheme. If G is a $k$-superscheme, we call it an algebraic supergroup.
Remark 3.15.
We notice, that since an algebraic supergroup is defined by limit diagrams (in this case pullbacks), that whenever we have a limit preserving functor (e.g. the inclusion or bosonic reduction as both of them are right adjoints), that group objects will be sent to group objects.
Another standard way to think about superschemes, is in terms of representable functor
Definition 3.16.
Let $\mathcal{C}$ be a locally small category and let $\operatorname{Set}$ be the category of sets. For each object $A$ of $\mathcal{C}$ let $\operatorname{Hom}(A,-)$ be the hom functor that maps object X to the set $\operatorname{Hom}(A,X)$.
A functor
(10) $\mathcal{F}:\mathcal{C}\rightarrow\operatorname{Set}$
is said to be representable if it is naturally isomorphic to Hom(A,–) for some object A of C. A representation of F is a pair $(A,\phi)$ where
(11) $\phi:\operatorname{Hom}(-,A)\rightarrow\mathcal{F}$
is a natural isomorphism.
Remark 3.17.
Affine schemes are representable by a functor from $\operatorname{Sch}^{\operatorname{op}}\rightarrow\operatorname{Set}$. Furthermore, in this setting one can think of an algebraic group as a functor from
$\mathcal{F}:\operatorname{sSch}^{\operatorname{op}}\rightarrow\operatorname{Grp},$
which we usually call a group functor.
Example 3.18 ([5]).
Let $V$ be a super vector space. Then the functor of points of the associated affine space is
[…]
If we now consider the $R$-points, where $R$ is a superalgebra. Consider a purely odd vector, it is of the form
(13) $v=\sum\alpha_{i}\otimes e_{i},$
so we were able to write down a vector, with odd coefficients, which can not be seen, by considering the pedestrian definition given earlier. This is important, because to prove a later result we have to work with $R$-points of matrices and vectors.
Example 3.19.
Consider $SL_{2}=\operatorname{Spec}(\mathbb{Z}[a,b,c,d]/(ad-bc-1))$. Then its group functor is given by
$\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[a,b,c,d]/(ad-bc-1),-)$
We can now compute
$\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[a,b,c,d]/(ad-bc-1),\mathbb{C})=SL_{2}(\mathbb{C})$
because we ring morphisms are determined by the evaluation maps for each variable. The evaluations live inside of $\mathbb{C}$ and have to fulfill the determinant condition. The result is precisely $\operatorname{SL}_{2}(\mathbb{C})$.
However, for now let us return to the first definition of a group superscheme. Assume $G$ to be an affine $k$-superscheme and apply the spec functor to this series of diagrams and we get the following commuting diagrams:
[…]
So we have a co-group structure on our $k$-superalgebra, which is then called a Hopf superalgebra.
Remark 3.20.
If we now live in the affine setting then it also turns out that we have a functor from
$\operatorname{SAlg}_{k}\rightarrow Grp.$
An easy consequence of this remark is, that for a $k$ Grassman algebra $\Lambda$ we have the natural quotient map from $\Lambda\rightarrow k$. We use the fact, that we indeed have a group functor which results in the induced morphism
$G(\Lambda)\rightarrow G(k)=G_{red}(k).$
Another important notion in the theory of algebraic supergroups, while not immediately relevant here, is the notion of a Harish-Chandra pair. We choose to write this short section as we believe it is quite helpful in general when thinking about algebraic supergroups. It turns out that this category is equivalent to the category algebraic supergroups [15] [9].
Definition 3.21.
Suppose $(G_{0},\mathfrak{g})$ are respectively a group (analytic or affine algebraic) and a super Lie algebra. Assume that:
  1. (1)
    $\mathfrak{g}_{0}\simeq\mathrm{Lie}(G_{0})$,
  2. (2)
    $G_{0}$ acts on $\mathfrak{g}$ and this action restricted to $\mathfrak{g}_{0}$ is the adjoint representation of $G_{0}$ on $\mathrm{Lie}(G_{0})$. Moreover the differential of such action is the Lie bracket. We shall denote such an action with $\mathrm{Ad}$ or as $g.X$, $g\in G_{0}$, $X\in\mathfrak{g}$.
Then $(G_{0},\mathfrak{g})$ is called a super Harish-Chandra pair (SHCP).
Let SHCP be the category of super Harish-Chandra pairs and ASP the category of algebraic supergroups then as mentioned earlier we also have the following equivalence of categories
(14) $\textbf{SHCP}\cong\textbf{ASP}$
due to Gavarini [9] and proven later again over commutative rings by Masuoka and Shibata in [15].
Remark 3.22.
In [9] the equivalence is essentially constructed by exponentiating all odd elements of the Lie superalgebra. This is well-defined, because due to the nilpotence of odd elements the exponential will truncate at first order. As an example consider $\mathbb{C}[\theta]$ with $\theta^{2}=0$ thus we have.
(15) $\operatorname{exp}(\theta)=1+\theta$
By the same token, we will also see that the Hopf algebra, that corresponds to the algebraic supergroup, can be decomposed as a superscheme into
(16) $k[G]=k[G_{0}]\otimes\bigwedge\mathfrak{g}_{1}^{\vee}$
Remark 3.23.
The group structure is not necessarily preserved under this isomorphism, however the co-unit is preserved [15].
The supergroups whose group structure is preserved under the splitting, were also already classified. It turns out it is equivalent to $\mathfrak{g}_{1},\mathfrak{g}_{1}]=0 [$24].
$3.$ $\operatorname{OSp}(1|2)$
Let $V\coloneqq\mathbb{R}^{1|2}$ be a super vector space over the Grassmann-algebra $\Lambda$ and $B$ be an even non-degenerate bilinear form.
Definition 3.24 ([14]).
Consider the vector space $V$ and $B$ as above, then
(17) $\operatorname{OSp}(1|2)\coloneqq\{g\in\operatorname{GL}(V)|B(gv,gw)=B(v,w)\ \forall v,w\in V\}$
The group law is given by ordinary matrix multiplication. Alternatively, we can also describe it as a group preserving a super symplectic structure $\omega$ on $\mathbb{R}^{2|1}$ [18], so there can arise confusion due to this $\operatorname{OSp}(1|2)$ acting $\mathbb{R}^{2|1}$. The descriptions are equivalent and for now we will proceed with the symplectic one. Generally we will be working over $\mathbb{C}$ and denote this by $\operatorname{OSp(1|2)}$. We can also think of $\operatorname{OSp}(1|2)$ as a group functor from the category of algebraic superschemes to the category of groups. We recover the previous notion of $\operatorname{OSp}(1|2)$ matrices via the functor of points and thinking of the $\Lambda$-points as matrices of the form
$\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}$
with $a,b,c,d,f\in\Lambda_{0}$ and $\alpha,\beta,\gamma,\delta\in\Lambda_{1}$. These matrices are also subject to some relations, but before we write them out, we should also know where they come from. Since we are defining a group preserving the symplectic form, we can also write out the symplectic form as a matrix which is
$J=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix}$
which should be invariant under the action of $\operatorname{OSp}(1|2)$. From this we also have a nice formula for the inverse of an element $g$. Let $g\in\operatorname{OSp}(1|2)$, then its inverse is given by
$g^{-1}=J^{-1}g^{st}J,$
which in formulas is
$g=\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\quad,\quad g^{-1}=\begin{pmatrix}d&-b&\delta\\ -c&a&-\gamma\\ -\beta&\alpha&f\end{pmatrix}$
By explicit calculation, one can also see that these properties force the relations
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc.$
onto us and we impose the relation $\operatorname{Ber}(g)=1$ separately. One can also notice, that the underlying bosonic group is $\operatorname{SL}(2,-)$. It follows that we naturally get a surjective ring homomorphism induced by quotienting all odd generators and we get a closed immersion of schemes
[…]
i.e.
(18) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\longmapsto\begin{pmatrix}a&b&0\\ c&d&0\\ 0&0&1\end{pmatrix}$
By the functorial nature of algebraic groups, we can now also consider the natural morphism from a $k$-Grassmann algebra $\Lambda$ to $k$, by sending all odd generators to 0. This induces the morphism
$\operatorname{OSp}_{1|2}(\Lambda)$$\operatorname{OSp}_{1|2}(k)=SL_{2}(k)$}
and the identification can be done with the functor of points in the same way as we did earlier for the functor $SL_{2}(-)$, so in formulas we have
(19) $\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}&0\\ c_{\#}&d_{\#}&0\\ 0&0&1\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}$
Furthermore, when we take a look at the composition
[…]
We get a splitting for the supergroup $\operatorname{OSp}_{1|2}$. We can once again introduce notions like semisimple or reductive, however being a semisimple algebraic supergroup is a very special phenomena. The classification of semisimple algebraic supergroups is well known and was originally proven by Hochschild and Djokovic [7]. Recently Sherman reproved this with more modern techniques in [21]. We will state Shermans version.
Theorem 3.25 ([21]).
Let $G$ be a connected algebraic supergroup. Then $Rep(G)$ is semisimple if and only if
(20) $G=K\times\operatorname{OSp}(1|2n_{1})\times...\times\operatorname{OSp}(1|2n_{k})$
for a reductive algebraic group $K$ and positive integers $n_{1},...n_{k}$.
Remark 3.26.
We can see that in the special case we will be tackling, we are still working with a reductive algebraic supergroup, which has nice properties with respect to taking quotients.
Next we describe the mentioned triangulation lemma for $\operatorname{OSp}(1|2)$.
Lemma 3.27.
Let $A_{0}\in SL(2,\mathbb{C})\subset\operatorname{OSp}(1|2)$ and $B_{0}\in\operatorname{OSp}(1|2)$. Furthermore, assume that $A$ can be diagonalized as $A=diag(\mu,\mu^{-1},1)$ and $\mu\neq\mu^{-1}$, then one can always find an element $g\in\operatorname{OSp}(1|2)$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
$gA_{0}g^{-1}=\begin{pmatrix}\mu&0&0\\ 1&\mu^{-1}&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\psi\\ \psi&0&1\end{pmatrix}\quad,\quad gB_{0}g^{-1}=\begin{pmatrix}\lambda&\kappa&0\\ 0&\lambda^{-1}&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&\xi\\ 0&1&0\\ 0&-\xi&1\end{pmatrix}$
Proof.
Let
$g=\begin{pmatrix}0&-x^{-1}&0\\ x&y&\nu\\ 0&-x^{-1}\nu&1\end{pmatrix}$,
where $x,y,\nu$ are indeterminates and conjugate $diag(\mu,\mu^{-1},1)$, which yields
$\begin{pmatrix}\mu^{-1}&0&0\\ xy(\mu-\mu^{-1})&\mu&\nu(-\mu+1)\\ \nu(\mu^{-1}-1)&0&1\end{pmatrix}$
define $\psi\coloneqq\nu(\mu^{-1}-1)$ and choose $xy=(\mu-\mu^{-1})^{-1}$ and we can reduce this to
$A=\begin{pmatrix}\mu^{-1}&0&0\\ 1&\mu&\psi\mu\\ \psi&0&1\end{pmatrix}$.
On the other hand we now have to conjugate the matrix $B_{0}$:
$CB_{0}C^{-1}=\begin{pmatrix}0&-x^{-1}&0\\ x&y&\nu\\ 0&-x^{-1}\nu&1\end{pmatrix}\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\begin{pmatrix}y&-x^{-1}&-x^{-1}\nu\\ x&0&0\\ -\nu&0&1\end{pmatrix}$
The result is rather non enlightening. However, we can use the same reasoning as in the $\operatorname{SL}(2,\mathbb{C})$ case and set the bosonic bottom left corner to be zero. So let us look at the argument and write
$B=CB_{0}C^{-1}=\begin{pmatrix}S&T&\epsilon\\ P&Q&\eta\\ S\eta-P\epsilon&T\eta-Q\epsilon&1+\epsilon\eta\end{pmatrix}$,
where $P=-bx^{2}+cy^{2}+axy-dxy+\alpha\nu+y\beta\nu$. Now set $P=0$, because as previously seen we did not make a choice for $y$ yet, so we can find a solution to that polynomial. This reduces our problem to
$B=\begin{pmatrix}S&T&\epsilon\\ 0&Q&\eta\\ S\eta&T\eta-Q\epsilon&1+\epsilon\eta\end{pmatrix}$,
where we notice, that $\eta$ is a function of $\beta$, so we work with the object $\eta(\nu)$. The function is naturally linear in $\nu$, so we can simply solve for $\nu$ and due to the upper triangular shape we also rewrite $S=\lambda$, $T=\kappa$ and $Q=\lambda^{-1}$ and we arrive at
$B=\begin{pmatrix}\lambda&\kappa&\epsilon\\ 0&\lambda^{-1}&0\\ 0&\lambda^{-1}\epsilon&1\end{pmatrix}$.
Finally, define $\xi\coloneqq\lambda^{-1}\epsilon$ can write down the normal form
$\begin{pmatrix}\lambda&\kappa&\lambda\xi\\ 0&\lambda^{-1}&0\\ 0&\xi&1\end{pmatrix}$.
This explicit computation, now lets us write down ”invariant” odd functions $\theta,\xi$ (invariant up to $\pm$). At this point we have to write down what $\nu$ is. In the conjugation one computes that
$\epsilon$ $=-x^{-1}\beta$
$\eta$ $=-x^{-1}(xa+yc)\beta+(x\alpha+y\nu+\beta f)$
setting $\eta=0$ one can solve for $\nu$ which yields
$\nu$ $=y^{-1}(x^{-1}(xa+yc)\beta-x\alpha-\beta f).$
Collecting all of the previous results we can write down
$\psi$ $=y^{-1}(x^{-1}(xa+yc)\beta-x\alpha-\beta f)(\mu^{-1}-1)$
$\xi$ $=-\lambda^{-1}x^{-1}\beta,$
where we remind ourselves that $x=y^{-1}(\lambda-\lambda^{-1})^{-1}$ and $y$ is a solution of the polynomial $P$. We are working in a non-integral ring, so the existence of a solution is non-obvious. However, a result of Shigeaki Nagamichi [12], tells us as long as the reduction has roots, the corresponding super polynomial also has roots. ∎
Remark 3.28.
Although it is proven for complex coefficients, we can see that the same proof strategy works for the case of $\operatorname{OSp}(1|2)$ with real coefficients, where the reduction of that element is hyperbolic.
Remark 3.29.
In Witten and Stanford [23] they make use of this parametrization of the moduli space of flat $\operatorname{OSp}(1|2)$-connections to compute an explicit form of $G$-invariant measures. They also write it down in the form
$U$ $=\begin{pmatrix}e^{a}&\kappa&0\\ 0&e^{-a}&0\\ 0&0&1\end{pmatrix}exp(\xi q_{1})$ $V=\begin{pmatrix}e^{-b}&0&0\\ 1&e^{b}&0\\ 0&0&1\end{pmatrix}exp(\xi q_{2})$
$q_{1}$ $=\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&-1&0\end{pmatrix}$ $q_{2}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$
This might look different compared to the parametrization given earlier, however one remembers that the Taylor expansion of odd things vanishes after the first order. Expanding it that way gives you the desired result. Furthermore, one notes that this parametrization of our pair $(A,B)$ ensures that there are 3 even degrees of freedom and 2 odd degrees of freedom (defined up to sign). We also note, that $q_{1},\ q_{2}$ are the generators of $\mathfrak{g}_{1}$ of $\mathfrak{osp}(1|2)$ which you simply exponentiate to turn them into elements in the corresponding algebraic supergroup.
3.4. Invariant Theory
In this section, we want to introduce the necessary tools from classical invariant theory and adapt them to the $\mathbb{Z}_{2}$-graded case. Most of these results will generalize without adapting the proof. We follow [13] and [17]
First a short remark on $\operatorname{OSp}(1|2)$. Since we have semisimple representations for $\operatorname{OSp}(1|2)$ we also get that it is linearly reductive. We start by showing the so-called Reynolds operator exists, which is important for the following decomposition.
Definition 3.30 ([17]).
Let $R:V\rightarrow V$ be a linear map of a vector space $V$. Let $G$ be an algebraic group G, which acts on V rationally. It is called a Reynolds operator if $R$ is a projection onto $V^{G}$, the space of invariant vectors in $V$.
Suppose this operator exists and let $Y$ be a subvariety of $X$. Then we have a short exact sequence of $G$-modules, which by linear reductivity gives us also a splitting and hence the diagram
[…]
The induced map on invariant subrings is surjective by the Reynolds property and we can check injectivity by diagram chasing.Therefore,
(21) $k[Y]^{G}=k[X]^{G}/I^{G}.$
Let us employ the outline strategy in a specific example, i.e. let us reprove the result of Fricke-Klein again.
Example 3.31.
Consider the ring $\mathbb{C}[M_{2}^{2}]^{SL_{2}}$, the ring of invariants of $2\times 2$ matrices invariant under conjugation by $SL_{2}$. It is known that this ring is
(22) $\mathbb{C}[M_{2}^{2}]^{SL_{2}}=\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB),\operatorname{tr}(A^{2}),\operatorname{tr}(B^{2})].$
the ideal $I$ is generated in this case by $I=\langle\operatorname{det}(A)-1,\operatorname{det}(B)-1\rangle$. At first one might think, that this will not quotient out any generators but we can remedy that by a simple application of the Cayley-Hamilton theorem. For a $2\times 2$ matrix $A$ it gives us the equation:
(23) $A^{2}-\operatorname{tr}(A)A+\operatorname{det}(A)\operatorname{id}=0,$
which after applying the trace to the whole equation gives us
(24) $\begin{split}\operatorname{tr}(A^{2})-\operatorname{tr}(A)^{2}+2=0\\ \operatorname{tr}(A^{2})=\operatorname{tr}(A)^{2}-2.\end{split}$
This now makes the generators $\operatorname{tr}(A^{2}),\operatorname{tr}(B^{2})$ redundant and we get
(25) $\mathbb{C}[M_{2}^{2}]^{SL_{2}}/I^{SL_{2}}=\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)].$
One should be careful though, that this isomorphism is an isomorpism of modules and not of rings. However, it does give us enough information about the nature and number of the generators which is enough for us.
Remark 3.32.
Even though this proof used the Cayley-Hamilton theorem, it is different to the proof given in [10].
Let us establish, that indeed for a linearly reductive supergroup, we can once again define the Reynolds operator.
Theorem 3.33.
Let $G$ be a linearly reductive algebraic supergroup that acts rationally on a $k$-vector space $V$ and let $V^{G}$ be the subspace of invariant vectors. Then there exists a uniquely defined linear operator
(26) $R:V\longrightarrow V$
projecting $V$ onto $V^{G}$.
Proof.
If $v\in V$ is not invariant, then it is contained in some subspace $W$, such that $G\cdot W\subset W$. By linear reductivity $W$ once again splits into
(27) $W=W^{G}\oplus W^{\prime}$
It follows that $v\in W^{\prime}$, so it is non-empty invariant and $W^{\prime}\cap V^{G}=0$. By applying Zorn’s lemma to the set of invariant subspaces, with this property we can find a maximal one, which we are gonna call $V^{\prime}$. Suppose, there are two such maximal subspaces, which we are going to denote by $V_{1},V_{2}$. It holds for the direct sum that $V_{1}\oplus V_{2}\cap V^{G}=0$. To show that, assume that it is non-empty, thus we can find an element $w=v_{1}\oplus v_{2}$, which is invariant under the group action. However, since the action acts on both parts separately, this is a contradiction, hence $V_{1}\oplus V_{2}\cap V^{G}=0$. This means, that the direct sum, would be a larger subspace with the wanted property, which once again is a contradiction. Thus $V^{\prime}$ is uniquely determined and we can also see that
(28) $V^{G}\oplus V^{\prime}=V.$
Now we define $R$ by the property, that it fixed $V^{G}$ and has $V^{\prime}$ as its kernel. ∎
Remark 3.34.
From here on out, we could prove Nagata’s theorem in the supercase, however that is not really useful, since previously we have seen that the class of redudctive supergroups is not very large. Therefore, it is not nearly as powerful as in the case of algebraic groups.
We continue by establishing that polarization and restitution works just as in the purely bosonic case. For the reader unfamiliar with classical invariant theory, polarization allows us to transform the polynomial into a multilinear expression, where we already know the invariants and restitution is the inverse process. We combine in the following way
PolarizationComputing InvariantsRestitution
which allows us to compute the polynomial invariants, which arise out of the multilinear ones. To be a little bit more precise, we want to show that
$\bigoplus_{d}k[V]_{d}$$\bigoplus_{d}k[V^{d}]$$\bigoplus k[V]_{d}$$\scriptstyle{\mathcal{P}}$$\scriptstyle{F}$
is $G$-equivariant, where the maps will be explained shortly. We can see that one can simply chase an invariant polynomial $f\in k[V]$ through this construction, find an explicit form in terms of multilinear expressions and then recover the original polynomial in terms of the multilinear invariants. This strategy shows, why it is enough to be able to compute multilinear invariants, which is much easier compared to the more difficult problem of finding polynomial invariants right off the bat.
Definition 3.35.
Consider a direct sum $V=V_{1}\oplus...\oplus V_{r}$ of finite dimensional super vector spaces. A function $f\in k[V_{1}\oplus...\oplus V_{r}]$ is called mukltihomogeneous of degree $h=(h_{1},...,h_{r})$ if $f$ is homogeneous of degree $h_{i}$ in $V_{i}$, i.e. for all $v_{1},...,v_{r}\in V$ and $t_{1},...,t_{r}\in k$ we have
(29) $f(t_{1}v_{1},...,t_{r}v_{r})=t_{1}^{h_{1}}\cdot...\cdot t_{r}^{h_{r}}f(v_{1},...v_{r}).$
One can also call $h$ the multidegree of $f$.
Every polynomial function $f$ is in a unique way a sum of multihomogeneous functions $f=\sum f_{h}$. The $f_{h}$ are usually called multihomogenous components of $f$. This gives rise to a decomposition
(30) $k[V_{1},...,V_{r}]=\bigoplus_{h\in\mathbb{N}^{r}}k[V_{1},...,V_{r}]_{h},$
where $k[V_{1},...,V_{r}]_{h}$ denotes the subspace of multihomogeneous functions of degree $h$.
Remark 3.36.
This is a different grading from the natural $\mathbb{Z}_{2}$-grading of our superring.
Remark 3.37.
The way this interacts with super is demonstrated in the following. Assume $V_{r}$ is entirely odd, then it means the polynomial $f$ can be at most homogeneous of degree $h_{r}=1$
It is clear that the multidegree is preserved under the action of $\operatorname{GL}(V_{1})\times...\times\operatorname{GL}(V_{r})$. Furthermore, if $V_{1},...V_{r}$ are representations, than we also have a natural grading on the ring of invariants
(31) $k[V_{1},...,V_{r}]^{G}=\bigoplus_{h\in\mathbb{N}^{r}}k[V_{1},...,V_{r}]_{h}^{G}.$
Definition 3.38.
Let $V$ be a finite dimensional $k$-vector space. Identify $k[V]=\bigoplus_{d}k[V]_{d}$, where $k[V]_{d}$ denotes the space of homogeneous polynomials of degree $d$. Let $f\in k[V]_{d}$ and $v=\sum_{i}^{d}t_{i}v_{i}$ for $v_{i}\in V$. We define
$\mathcal{P}(f)(\sum_{i}t_{i}v_{i})=\sum_{s_{1}+...+s_{d}=d}t_{1}^{s_{1}}...t_{d}^{s_{d}}f_{s_{1},...,s_{d}}(v_{1},...,v_{d}),$
where $f_{s_{1},...,s_{d}}\in k[V^{d}]$ are well-defined and multihomogeneous of multidegree $(s_{1},...,s_{d})$. We call $f_{1,...,1}\in k[V^{d}]$ the (full) polarization of $f$ denoted by $\mathcal{P}f$
Lemma 3.39.
The linear operator $\mathcal{P}\colon K[V]_{d}\to K[V^{d}]_{(1,1,\ldots,1)}$ has the following properties:
  1. (a)
    $\mathcal{P}$ is $\mathrm{GL}(V)$-equivariant;
  2. (b)
    $\mathcal{P}f$ is symmetric;
  3. (c)
    $\mathcal{P}f(v,v,\ldots,v)=d!\,f(v)$.
Proof.
For equivariance consider the diagram
$k[V]_{d}$$k[V^{d}]$$k[V]_{d}$$k[V^{d}]$$\scriptstyle{\mathcal{P}}$$\scriptstyle{g\cdot}$$\scriptstyle{g\cdot}$$\scriptstyle{\mathcal{P}}$
and let
(32) $f(t_{1}v_{1}+...+t_{d}v_{d})=\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(v_{1},...v_{d}).$
We start chasing
$\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(v_{1},...v_{d})$$t_{1}\cdot...\cdot t_{d}f_{s_{1}...s_{d}}(v_{1},...v_{d})$$\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(g^{-1}v_{1},...,g^{-1}v_{d})$$t_{d}f_{s_{1}...s_{d}}(g^{-1}v_{1},...g^{-1}v_{d})$$\scriptstyle{g\cdot}$$\scriptstyle{g\cdot}$
$\mathcal{P}$ is also obviously symmetric. It remains to show $(c)$. Let us calculate
(33) $f(t_{1}v,...,t_{d}v)=(\sum_{i}t_{i})^{d}f(v)=(d!t_{1}\cdot...\cdot t_{d}+...)f(v).$
Looking at the first factor the claim follows. ∎
Definition 3.40.
For a multilinear $F\in K[V^{d}]$ the homogeneous polynomial $\mathcal{R}F$ defined by $\mathcal{R}F(v):=F(v,v,\ldots,v)$ is called the (full) restitution of $F$.
By construction we also have the following result.
Proposition 3.41.
Assume $\operatorname{char}k=0$ and let $V$ be a finite dimensional representation of a supergroup $G$. Then every homogeneous invariant $f\in K[V]^{G}$ of degree $d$ is the full restitution of a multilinear invariant $F\in K[V^{d}]^{G}$.
Proof.
Again, $\mathcal{R}\colon K[V^{d}]_{(1,\ldots,1)}\to K[V]_{d}$ is a linear $\mathrm{GL}(V)$-equivariant operator and we have $\mathcal{R}\mathcal{P}f=d!f$ by the property (c) of the lemma above. In fact, $f$ is the full restitution of $\frac{1}{d!}\mathcal{P}f$, which is a multihomogeneous invariant. ∎
We continue building up all necessary results, now we will begin describing how to translate vector invariants to matrix invariants. Firstly, we will state the first fundamental theorem for $GL(V)$
Theorem 3.42.
Let $V$ be a finite dimensional super vector space. The multilinear invariants under the natural $GL(V)$-action are given by
(34) $\begin{split}V^{\otimes N}\otimes(V^{\vee})^{\otimes N}&\longrightarrow\mathbb{C}\\ v_{1}\otimes...\otimes v_{N}\otimes\phi_{1}\otimes...\otimes\phi_{N}&\longmapsto\prod_{i}\langle v_{i},\phi_{\sigma(i)}\rangle\end{split}$
and all its transforms under the symmetric group
It remains to show, that the pairings are still given by traces, which is essentially just a retelling of Procesi [20], but with the word super attached.
Proposition 3.43.
Let V be a $p|q$ dimensional super vector space and $V^{\vee}$ its dual. We take $\langle\cdot,\cdot\rangle:V\otimes V^{\vee}\rightarrow k$, then
$\langle v,\phi\rangle=\operatorname{str}(v\otimes\phi)$
Proof.
This follows by a simple calculation. Take a homogeneous basis $\left\{e_{j}\right\}$ for $V$ and let $\left\{\varphi_{j}\right\}$ be its dual basis. We take a vector $v\otimes\varphi\in V\otimes V^{\vee}$ and compute the coefficients of the associated matrix $A$. Let
$v$ $=\sum_{i}a_{i}e_{i}\quad,\ \text{such that}\ \forall|e_{i}|=\{0,1\}:\ a_{i}\in\Lambda_{\{0,1\}}$
$\varphi$ $=\sum_{j}b_{j}\varphi_{j}\quad,\ \text{such that}\ \forall|\varphi_{j}|=\{0,1\}:\ b_{j}\in\Lambda_{\{0,1\}}$
and we compute
$v\cdot\varphi=\sum_{ij}a_{i}b_{j}e_{i}\otimes\varphi_{j}$
on the other hand we can also compute
$\langle v,\varphi\rangle=\varphi\cdot v=\sum_{i}(-1)^{|e_{i}||\varphi_{i}|}a_{i}b_{i}.$
It is easy to see that
$\operatorname{str}(A)=\langle v,\varphi\rangle$
Lemma 3.44.
Let $V$ be a super vector space. Then the space multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is isomorphic to the space of multilinear functions of $\operatorname{End}(V)^{m}$
Proof.
The space of multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is given by
(35) $(V^{\otimes m}\otimes(V^{\vee})^{\otimes m})^{\vee}\cong((V\otimes V^{\vee})^{\otimes m})^{\vee}\cong(\operatorname{End}(V)^{\otimes m})^{\vee}.$
We immediately recognize, that the right hand side are the multilinear functions on $\operatorname{End}(V)^{m}$. ∎
Remark 3.45.
All these spaces naturally admit a $\operatorname{GL}(V)$-action. Considering all these vector spaces as $\operatorname{GL}(V)$-representations, it is also easy to see, that these identifications are a $GL(V)$-equivariant isomorphism.
We can now translate the multilinear invariant in vector form into the multilinear invariant in matrix form. Consider $\operatorname{End}(V)^{\otimes i}$ and choose $\sigma\in Sym_{i}$. Decompose $\sigma$ into disjoint cycles
(36) $\sigma=(i_{1}i_{2}...i_{k})(j_{1}j_{2}...j_{n})...(t_{1}...t_{e})$
including the ones of length 1.
Theorem 3.46.
Given $A_{1},...,A_{i}\in\operatorname{End}(V)$, we have:
(37) $\mu_{\sigma}(A_{1}\otimes...\otimes A_{i})=\operatorname{str}(A_{i_{1}}...A_{i_{k}})\operatorname{str}(A_{j_{1}}...A_{j_{n}})\operatorname{str}(A_{t_{1}}...A_{t_{e}})$
Proof.
Since both sides are multilinear, it is enough to prove it for the case where $A_{i}=\phi_{i}\otimes v_{i}$.
(38) $\mu_{\sigma}(A_{1}\otimes...\otimes A_{i})$ $=\prod_{i}\langle\phi_{\sigma(i)},v_{i}\rangle$
(39) $=\langle\phi_{i_{2}},v_{i_{1}}\rangle\langle\phi_{i_{3}},v_{i_{2}}\rangle...\langle\phi_{t_{1}},v_{t_{e}}\rangle$
Consider, for instance, the product
(40) $M$ $=\langle\varphi_{i_{2}},v_{i_{1}}\rangle\langle\varphi_{i_{3}},v_{i_{2}}\rangle\cdots\langle\varphi_{i_{k}},v_{i_{k-1}}\rangle\langle\varphi_{i_{1}},v_{i_{k}}\rangle.$
One verifies immediately that
(41) $\varphi_{i_{1}}\otimes v_{i_{1}}\cdot\varphi_{i_{2}}\otimes v_{i_{2}},\ldots,\varphi_{i_{k}}\otimes v_{i_{k}}$ $=\langle\varphi_{i_{2}},v_{i_{1}}\rangle\langle\varphi_{i_{3}},v_{i_{2}}\rangle\cdots\langle\varphi_{i_{k}},v_{i_{k-1}}\rangle\varphi_{i_{1}}\otimes v_{i_{k}}.$
Therefore,
(42) $M=\mathrm{tr}(A_{i_{1}}A_{i_{2}}\cdots A_{i_{k}}),$
The last bit of prep work we have to do now is simply state the first fundamental theorem of invariants for $\operatorname{OSp}(1|2)$. Let $V$ be a finite dimensional super vector space with $dim(V)=m|2n$ and let $B$ be the symmetric pairing preserved under $\operatorname{OSp}(V)$. Deligne, Lehrer and Zhang proved a first fundamental theorem for $\operatorname{OSp}(V)$ [5].
Theorem 3.47.
Let $V$ be as above, then the multilinear $\operatorname{OSp}(V)$-invariant functions on $V^{N}$ for $N$ even are given by
(43) $\begin{split}V^{\otimes N}&\longrightarrow\mathbb{C}\\ v_{1}\otimes...\otimes v_{N}&\longmapsto B(v_{1},v_{2})\cdot...\cdot B(v_{N-1},v_{N}),\end{split}$
and all its permutations in $Sym_{N}$.
This was also proven by Lehrer and Zhang in [14] with a different method of proof earlier than the stated result mostly in the language of Brauer diagrams. With that method they also computed the polynomial invariants ones from the multilinear ones, which are simply the monomials in the multilinear invariants.
Theorem 3.48.
Let $V$ be as above, then the polynomial $\operatorname{OSp}(V)$-invariants are polynomials in the variables $B(v_{i},v_{j})\coloneqq(i,j)$ $\forall 1\leq i<j\leq m$
We can now prove a second fundamental theorem for the polynomial invariants. to do that, we need to prove a small lemma first.
Lemma 3.49.
Let $G\coloneqq(B(v_{i},v_{j}))_{i,j}$ then $\operatorname{det}(G)\neq 0$ iff $v_{1},...v_{N}$ are linearly independent.
Proof.
We proof this by contrapositive. If $\operatorname{det}(G)=0$ then there exists a relation between the rows $R_{i}\coloneqq(B(v_{i},v_{1}),...,B(v_{i},v_{N}))$, i.e.
(44) $\sum_{i}a_{i}R_{i}=0$
for some non-zero $a_{i}$. It follows
(45) $\sum_{i}a_{i}R_{i}=(B(\Sigma_{i}v_{i},v_{1}),...,B(\Sigma_{i}v_{i},v_{N}))=0$
So each factor must be zero, thus consider the linear combination
(46) $\begin{split}a_{1}B(\Sigma_{i}v_{i},v_{1})+...+a_{N}B(\Sigma_{i}v_{i},v_{N})&=0\\ B(\Sigma_{i}v_{i},\Sigma_{i}v_{i})&=0\end{split}$
by non-degeneracy, they are linearly dependent. For the other direction, we also consider the contrapositive. if $v_{1},...,v_{N}$ are linearly dependent. Let $a_{i}$ be the coefficient such that the linear combination of the vectors vanishes, then consider the sum of the rows
(47) $\begin{split}&=\sum_{i}a_{i}R_{i}\\ &=(B(\Sigma a_{i}v_{i},v_{1}),...,B(\Sigma a_{i}v_{i},v_{N}))\\ &=B(B(0,v_{1}),...,B(0,v_{N}))=0\end{split}$
so there is a relation between the rows, thus the determinant must vanish. ∎
Theorem 3.50 (SFT).
Let everything be as above, then the polynomial ring of $\operatorname{OSp}(V)$-invariants is
(48) $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}/I_{n}$
Proof.
We will be investigating the kernel of the morphism of rings
(49) $\phi:\mathbb{C}[(i,j)]\longrightarrow\mathbb{C}[V^{m}]^{\operatorname{OSp(V)}}.$
Notice, that we can fit all the inner products into a matrix $G\coloneqq(B(v_{i},v_{j}))_{i,j}$. One also sees that $G\in\operatorname{Mat}_{N\times N}(\Lambda)$. By the earlier lemma we follow that the maximal rank of this matrix is $n+2m$. It follows, that all $(n+2m+1)\times(n+2m+1)$ minors must vanish, thus the morphism factors in the following way
$\mathbb{C}[(i,j)]$$\mathbb{C}[V^{m}]^{\operatorname{OSp(V)}}$$\mathbb{C}[(i,j)]/I_{n}$}
It is easy, to see that $I_{n}\subset\operatorname{ker}(\phi)$. Furthermore, let $f\in\operatorname{ker}(\phi)$. We use the previous factoring
[…]
where $I_{n}$ is the ideal generated by all $(n+1)\times(n+1)$ minors. we need to check that $\operatorname{ker}(\phi)\subset I_{n}$, which means we have to determine
(50) $f\operatorname{mod}I_{n}\longrightarrow 0.$
We know that $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}$ is also generated by all $(i,j)$, thus the morphism is injective, hence $f\operatorname{mod}I_{n}=0$, therefore $\operatorname{ker}(\phi)=I_{n}$
(51) $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}=\mathbb{C}[(i,j)]/I_{n}$
Now as a corollary we have
Corollary 3.51.
The only relation for the invariant ring of $\mathbb{C}[V^{4}]^{\operatorname{OSp}(1|2)}$ is $\operatorname{det}(G)=0$
Proof.
for $\operatorname{OSp}(1|2)$ and for copies of $V$ that all $4\times 4$ minors must vanish, which means the determinant vanishing is the only relation. ∎
A consequence of this, is that we have 9 independent generators.
4. Character Varieties
We follow [22]. Let $G$ be a reductive algebraic group over k. When looking at the set $Hom(\Gamma,G)$, we can associate to it a ring $\mathcal{R}(\Gamma,G)$, and the $k$-points of $\operatorname{|}{Spec}(\mathcal{R}(\Gamma,G))|=\operatorname{Hom}(\Gamma,G)$. Since every finitely generated group can be constructed by imposing a set of relation on the free group $F_{n}$, we will start by constructing the representation variety of $F_{n}$.
Proposition 4.1.
Let $G$ be an algebraic group over $k$ and $F_{n}$ the free group on $n$ letters then coordinate ring of the representation variety is $\mathcal{R}(F_{n},G)=k[G]^{\otimes n}$.
Proof.
It is clear, that a morphism of groups between $F_{n}$ and $G$ is determined by the images of the generators, there are no relations and thus a representation of $F_{n}$ is simply a tuple of $n$ elements of $G$, hence a point in the representation variety is given by
(52) $\mathcal{R}(F_{n},G)=k[G]^{\otimes n}$
If we now impose relations on the free group, we get
$\operatorname{Hom}(\Gamma,G)=\operatorname{Spec}(k[G\times...\times G]/I_{\Gamma}).$
The ideal $I_{\Gamma}$ is generated by the relations induced by $\Gamma$, which tell you that the images of the generators still have to obey the relations of $\Gamma$, because otherwise it would not be a group homomorphism. By the quotient nature of $\mathcal{R}(\Gamma,G)$ it is not obvious if it is an integral domain, which over $\mathbb{C}$ allows you to identify Specm and Spec. This space algebraically classifies all so-called $G$-representations. However we do want to quotient out all conjugacy equivalent representations. Thus we define
Definition 4.2.
Let $G$ be a reductive algebraic group over $k$ and $\Gamma$ a finitely generated group, then we call the categorical quotient
(53) $\operatorname{Spec}(\mathcal{R}(\Gamma,G)^{G})$
the character variety.
Remark 4.3.
If we now only consider reductive algebraic groups then it turns out that $\operatorname{Specm}(\mathcal{R}(\Gamma,G)^{G})=\operatorname{Hom}(\Gamma,G)$ is a good categorical quotient and by Nagata’s theorem it is finitely generated. One should remark, that non-reduced points may still exists. Let $\operatorname{Hom}^{i}(\Gamma,G)$ be the subset of irreducible representations. Then we also have that
$\operatorname{Hom}^{i}(\Gamma,G)//G$
is a good geometric quotient, i.e. the set corresponds to the set theoretic quotient, where if we now consider fibers of the quotient map, then we will recover the orbits of $G$. In general, there should be no expectation, that the full character variety is actually a good geometric quotient, i.e. it is not the orbit space of the representation variety.
Recalling the theorem of Fricke and Klein, we can now reinterpret it as a special case of the character variety of the free group and in particular an explicit solution to the question of finding minimal generators of a character variety.
We can easily see, that there are no obstructions to replacing the algebraic group with an algebraic supergroup. The character variety of $\operatorname{OSp}_{\mathbb{R}}(1|2)$ has already been investigated from the analytic point of view for example in [19], where Penner-coordinates were generalized. It has also appeared in [23], where they gave an analogous recursion formula to Mirzakhani for super hyperbolic surfaces with heuristics coming from supergeometry. The statement was proven with algebro-geometric methods in [18].
Remark 4.4.
Seeing how this construction actually only cares about the Hopf algebras, we can also generalize this question to quantum groups. This will definitely work if we only consider group relations, that can be manipulated in a way such that no inverses appear, because if we consider the quantum group via algebra-valued matrices, then the inverse does not necessarily live in the same algebra. A heuristic construction was already given in [1]. We note, that for quantum groups one reasons purely in the dual picture of the Hopf algebra, so one has to think about coinvariants of the conjugation coaction.
5. Main Result
Let us return from the geometric world and consider purely algebraic world again, where we now state our main result. Let $V$ be a $1|2$-dimensional super vector space.
Theorem 5.1.
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ has 7 independent generators.
Proof.
Let us consider the algebra
(54) $\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}.$
By polarization, we can carry the polynomial invariants to the multilinear invariants. Now we remember that
(55) $((\operatorname{End}(V)\otimes\operatorname{End})^{\vee})^{\operatorname{OSp}(V)}\cong((V\otimes V\otimes V\otimes V)^{\vee})^{\operatorname{OSp}(V)}.$
By the first fundamental theorem of the $\operatorname{OSp}(V)$ we know that the number of polynomial invariants of $V^{\otimes 4}$ is 10. By the second fundamental theorem of $\operatorname{OSp}(V)$ we know that there is one relation for $\operatorname{dim}(V)=1|2$ and thus 9 independent generators. Because the multilinear invariants of matrices and vectors are the same, the polynomial invariants are also the same. By reductivity we have
[…]
where $I$ is the defining ideal of $\operatorname{OSp}(V)^{2}$ as subvariety of $\operatorname{End}(V)^{2}$. It is generated by the relations
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc,$
together with $\operatorname{Ber}(A)=1$. We know that there exists no odd invariants on $\mathbb{C}[\operatorname{End}(V)^{2}]$ and we also know the co-action on $I$ sends the odd generators to odd generators. Thus all the odd relations, do not give us an invariant of $I$. However, the relation $\operatorname{Ber}(A)=1$ is invariant under the co-action, thus it is an invariant vector in $I$. It follows that
(56) $I^{\operatorname{OSp}(1|2)}=\langle\operatorname{Ber}(A),\operatorname{Ber}(B)\rangle.$
Therefore we have
(57) $\begin{split}\#\text{gen. }\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}&=\#\text{gen. }\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)}+\#\text{gen. }I^{\operatorname{OSp}(V)}\\ 9&=x+2\\ x&=7\end{split}$
By an earlier theorem, we know that all multilinear invariants are given by traces, thus all these invariants are traces as well. ∎
This result can be reinterpreted as the fact that the character variety of the free group on 2 letters of $\operatorname{OSp}(V)$ has 7 independent generators. Which is a deviation of the result for $SL(2,\mathbb{C})$.
There have been results, where in the real analytic case one can prove, that this space is actually of the expected real dimension $6|4$. One can easily see, why this fails in the algebraic case. We have seen that all invariants on the space of matrices, are given by a supertrace. We recall that
(58) $\operatorname{str}:\operatorname{OSp}(V)\longrightarrow\mathbb{C}^{1|0}$
and thus we cannot determine any odd generators with the help of traces. We can also interpret as an incarnation of the fact, that there is a natural $\mathbb{Z}_{2}$-action acting on the odd coordinates. Explicitly, it is
$=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&\xi\\ 0&1&0\\ 0&-\xi&1\end{pmatrix}\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$
$=\begin{pmatrix}1&0&-\xi\\ 0&1&0\\ 0&\xi&1\end{pmatrix}$
So this tells us, that if we take the quotient by the full group $\operatorname{OSp}(1|2)$, that the space has to be of purely bosonic dimension. One also sees by this calculation, that taking the full ring of invariants can only give you a non-reduced ordinary scheme, since the odd degrees of freedom are not invariant, as we always have to pair two odd coordinates, that will be flipped simultaneously to become an invariant.
This means what is $\operatorname{OSp}(1|2)$-invariant is the product of both odd guys. One reckons that one should consider the quotient $\operatorname{{POSp}}(1|2)\coloneqq\operatorname{OSp}p(1|2)/\mathbb{Z}_{2}$ instead.
Making this change, one also needs to consider the character variety of $POSp(1|2)$ instead.
Remark 5.2.
One might be disappointed by the structure of this character variety. However, as it turns out there are also other instances of moduli spaces of superobjects, that are of purely even nature like the picard scheme of a SUSY curve [3] or the coarse moduli of a superstack [4].