### v2

# A Super Version of a theorem of Fricke-Klein

Abstract.
In this paper, we study the character variety of the supergroup $\operatorname{OSp}(1|2)$ in the specific case of the free group on two letters as a first step towards an algebraic treatment for the character variety of supergroups. One can interpret this as an analogue of a result of Fricke and Klein, who studied this in the case of $\operatorname{SL}(2,\mathbb{C})$. We give a count of independent generators of the ring of invariants and discuss a few geometric consequences at the end.
1. Introduction
A result by Fricke and Klein [10] states, that the $SL(2,\mathbb{C})$-invariant functions on $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ are given by the ring
(1) $\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)].$
In the language of geometric invariant theory, we can identify the categorical quotient of the product variety $\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})$ by conjugation as affine space, i.e.
$\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})$$\operatorname{Spec}(\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)])\cong\mathbb{C}^{3}.$$\scriptstyle{\pi}$
Another geometric reinterpreation of this result is that the character variety of the free group on two letters and $SL(2,\mathbb{C})$ is isomorphic to $\mathbb{C}^{3}$, i.e.
(2) $\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}.$
In this paper we establish a variant of the theorem of Fricke and Klein for a supergeometric extension $\mathrm{SL}_{2}(\mathbb{C})$, specifically the complex supergroup $\mathrm{OSp}(1|2)$. The motivation of this result comes from the fact, that $\operatorname{OSp}(1|2)$ is the super generalization of $\operatorname{SL}(2,\mathbb{C})$ and takes a similar place in the theory of super Riemann surfaces, as $\operatorname{SL}(2,\mathbb{C})$ does in the classical theory. On the real analytic side of this story, the character variety is related to the super Teichmüller space, where each connected component of the space is homeomorphic to $\mathbb{R}^{6g-6|4g-4}/\mathbb{Z}_{2}$ [3] [25], where $\mathbb{Z}_{2}$ only acts on the odd coordinates.
We will show that the character variety for the free group on two letters $F_{2}$ and the supergroup $\mathrm{OSp}(1|2)$ fails to be a superscheme. Many results in supergeometry can be obtained by the same proof as in the classical setting and the literature frequently leaves the explicit verification to the reader. In this work we opt for a self-contained exposition for the necessary results adapted to the super setting. Organization of the paper. In section 2 we will give a short exposition on the result of Fricke and Klein. In section 3, we start by introducing basic notions of supergeometry, algebraic supergroups and invariant theory to understand our main result. In section 4 we prove our main result. In section 5 we introduce the character variety and the character stack, which is the geometric interpretation of our result, and discuss the geometric implication of the main result.
Theorem 1.1.
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ are generated by 7 independent traces.
Acknowledgments. I would like to thank my advisor Paul Norbury for his constant guidance and confidence in my work. I would also like to thank Gustav Berth, Miles Koumouris and Lior Yanovski for useful discussion.
I am supported by the University of Melbourne Graduate Research Scholarship.
2. Fricke-Klein
In this section, we give an elementary proof of the result of Fricke and Klein [10] and separately Vogt [33]. In more recent times, Goldman reproved this theorem in [12], where he used the Cayley-Hamilton theorem in an essential way in the proof. In this version, we forego the usage of the Cayley-Hamilton theorem and instead make use of a simultaneous conjugation of a pair of matrices into upper and lower triangular forms. Before we start proving the result of Fricke and Klein, let us first state the aforementioned triangulation lemma.
Lemma 2.1.
Let $A_{0}\in\operatorname{SL}(2,\mathbb{C})$ and $B_{0}\in\operatorname{SL}(2,\mathbb{C})$. Then one can always find an element $g\in\operatorname{SL}(2,\mathbb{C})$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
(3) $A_{0}\coloneqq gAg^{-1}=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}\quad,\quad B_{0}\coloneqq gBg^{-1}=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}$
Proof.
This splits into two cases. $A$ diagonalizable or not. If $A$ is diagonalizable, we assume that $A=\mathrm{diag(\mu,\mu^{-1})}$, such that $\mu\neq\mu^{-1}$. Consider
$C=\begin{pmatrix}0&-x^{-1}\\ x&y\\ \end{pmatrix}$
and compute its action by conjugation on A and B
$CAC^{-1}=\begin{pmatrix}\mu^{-1}&0\\ xy(\mu-\mu^{-1})&\mu\end{pmatrix}\quad,\quad CBC^{-1}=\begin{pmatrix}-x^{-1}yc+d&x^{-2}c\\ xya+y^{2}c-x^{2}b-xyd&a+x^{-1}yc\end{pmatrix}$
To set the bottom left corner of the first matrix to 1, we want $x=y^{-1}(\mu-\mu^{-1})^{-1}$. By substituting the $x$, the bottom left corner is equal to
$cy^{2}+\frac{a-d}{\mu-\mu^{-1}}-\frac{b}{\mu-\mu^{-1}}y^{-2},$
which we want to set to zero. So we can solve a polynomial of degree 4, to get an upper triangular matrix for $B$ and a lower triangular matrix $A$, by simultaneous conjugation.
If $A$ not diagonalizable, then it must admit Jordan normal form
(4) $A=\begin{pmatrix}\mu&1\\ 0&\mu\end{pmatrix}$.
As $A\in\operatorname{SL}(2,\mathbb{C})$, we have $\mu=\pm 1$, which is a case we can treat by hand. Consider
$\begin{pmatrix}x&y\\ 0&x^{-1}\end{pmatrix}\begin{pmatrix}\pm 1&1\\ 0&\pm 1\end{pmatrix}\begin{pmatrix}x^{-1}&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}\pm 1&x^{2}\\ 0&\pm 1\end{pmatrix}$
On the other hand for the arbitrary matrix we get
$\begin{pmatrix}x&y\\ 0&z\end{pmatrix}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\begin{pmatrix}z&-y\\ 0&x\end{pmatrix}=\begin{pmatrix}z(xa+yc)&-y(xa+yc)+x(xb+yd)\\ z^{2}c&-yzc+xzd\end{pmatrix}$
Also note that the reflection matrix acts on upper/lower triangular matrices as
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}a&b\\ 0&d\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}=\begin{pmatrix}a&0\\ -b&d\end{pmatrix}$
and
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}a&0\\ c&d\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}=\begin{pmatrix}a&-c\\ 0&d\end{pmatrix}$
First, we set $x=i$, then we choose a solution for the equation y(xa+yc)+x(xb+yd)-y(xa+yc)+x(xb+yd), which we can do over $\mathbb{C}$. We multiply by rotation matrix, which corresponds to rotating by $\frac{\pi}{2}$. Thus, we obtain the pair
$A=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\quad,\quad B=\begin{pmatrix}z(xa+yc)&-z^{2}c\\ 0&-yzc+xzd\end{pmatrix}$,
which is in the desired form. ∎
Remark 2.2.
There exists an analogous result for $\operatorname{OSp}(1|2)$, which will be proven later.
We are now ready to give a simple proof of the theorem mentioned in the introduction.
Theorem 2.3.
Let $A,B\in\operatorname{SL}(2,\mathbb{C})$ then all conjugation-invariant polynomial functions $f:SL(2,\mathbb{C})\times SL(2,\mathbb{C})\rightarrow\mathbb{C}$ are generated by $\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)$.
Proof.
Consider the coordinate ring of $SL(2,\mathbb{C})\times SL(2,\mathbb{C})$ denoted by $\mathbb{C}[G\times G]$ and its subring of $G$-invariant polynomial functions, denoted by $\mathbb{C}[H]$. We want to prove
[…]
where $X=\operatorname{tr}(A)$, $Y=\operatorname{tr}(B)$ and $Z=\operatorname{tr}(AB)$. First of all note that if $f\in\mathbb{C}[H]$, then $f$ is a polynomial function in the entries of the two matrices $(\xi,\eta)\in SL(2,\mathbb{C})\times SL(2,\mathbb{C})$. We will denote it $f(A,B)$. Furthermore, by the triangulation lemma we can bring them into the form
$A=\begin{pmatrix}\lambda&\kappa\\ 0&\lambda^{-1}\end{pmatrix}\quad,\quad B=\begin{pmatrix}\mu&0\\ 1&\mu^{-1}\end{pmatrix}$
This parametrization means, that we can view $f\in\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$, which makes $\mathbb{C}[H]$ into a subring by the obvious inclusion map
$\mathbb{C}[H]\rightarrow\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa].$
We can also define the map
$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$ $\rightarrow\mathbb{C}[X,Y,Z,\Delta_{X},\Delta_{Y}]$
$\lambda$ $\mapsto\frac{X+\Delta_{X}}{2}$
$\lambda^{-1}$ $\mapsto\frac{X-\Delta_{X}}{2}$
$\mu$ $\mapsto\frac{Y+\Delta_{Y}}{2}$
$\mu^{-1}$ $\mapsto\frac{Y-\Delta_{Y}}{2}$
$\kappa$ $\mapsto Z-\frac{X+\Delta_{X}}{2}\frac{Y+\Delta_{Y}}{2}-\frac{X-\Delta_{X}}{2}\frac{Y-\Delta_{Y}}{2},$
where $\Delta_{X}=\sqrt{X^{2}-4}$ and $\Delta_{Y}=\sqrt{Y^{2}-4}$. We can easily see that
[…]
If we now consider a polynomial $f\in\mathbb{C}[X,Y,Z]\subset\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$, it is easy to see it is $G$-invariant, thus $f\in\mathbb{C}[H]$. Notice that
$\lambda-\lambda^{-1}$ $\mapsto\Delta_{X}$
$\mu-\mu^{-1}$ $\mapsto\Delta_{Y}.$
We observe that the statement is even stronger, we have
[…]
Now we define a map $\psi$, by the composition of the solid arrows in the following diagram
$\mathbb{C}[H]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\mathbb{C}[X,Y,Z]$$\mathbb{C}[X,Y,Z,\Delta_{X},\Delta_{Y}]$$\scriptstyle{\eqqcolon\psi}$
which is surjective as argued earlier. We now prove it is injective. We again make use of our identifications
$\mathbb{C}[H]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\mathbb{C}[X,Y,Z]$$\mathbb{C}[\lambda,\lambda^{-1},\mu,\mu^{-1},\kappa]$$\scriptstyle{\operatorname{id}}$
Now consider the image of a $G$-invariant polynomial $g$. We have $\psi(g(\xi,\eta))=f(X,Y,Z)=0$. This is equivalent to
$\sum_{ijk}\alpha_{ijk}X^{i}Y^{j}Z^{k}=\sum_{ijk}\alpha_{ijk}(\lambda+\lambda^{-1})^{i}(\mu+\mu^{-1})^{j}(\lambda\mu+\lambda^{-1}\mu^{-1}+\kappa)^{k}=0,$
which we expand in the form
$\sum_{ijk}\alpha_{ijk}\sum_{l}^{i}\sum_{m}^{j}\sum_{k_{1}+k_{2}+k_{3}=n}^{k}\binom{i}{l}\binom{j}{m}\binom{k}{k_{1},k_{2},k_{3}}\lambda^{i-2l+k_{1}-k_{2}}\mu^{j-2m+k_{1}-k_{2}}\kappa^{k_{3}}$
On the other hand, we have the $G$-invariant polynomial
$\sum_{ijkpq}\beta_{ijkpq}\lambda^{i-j}\mu^{k-p}\kappa^{q}=0,$
which after relabeling turns into
$\sum_{rst}\beta_{rst}\lambda^{r}\mu^{s}\kappa^{t}=0.$
We observe that every $\beta_{rst}$ is a sum of the $\alpha_{ijk}$ By our condition, we know that all coefficients $\alpha_{ijk}$ must vanish, since there is no relation between $X,Y,Z$. Therefore $\beta_{rst}=0$. Thus $g=0$ and injectivity follows. In total we now have
$\mathbb{C}[G\times G]^{G}=\mathbb{C}[H]\cong\mathbb{C}[X,Y,Z],$
and on the level of spaces we have
$\operatorname{Hom}(F_{2},SL(2,\mathbb{C}))//SL(2,\mathbb{C})\cong\mathbb{C}^{3}.$
3. Preliminaries
3.1. Supergeometry
Here we want to introduce basic notions of supergeometry. In particular, we will provide a categorical viewpoint, alongside its classical definitions, thereby highlighting the similarities between commutative algebras and supercommutative algebra. This should aid the reader to understand which phenomena will directly carry over and in which situations some more care ought to be given. A possible way to view the category of commutative rings is that it is the category of commutative monoid objects with respect to the tensor product in the symmetric monoidal category $\mathrm{Ab}$. To do supercommutative algebra and eventually supergeometry, the only shift that will occur from the categorical perspective is that we replace the category $\mathrm{Ab}$ with $\mathrm{Ab}_{\mathbb{Z}_{2}}$. In practice, all objects will be endowed with a $\mathbb{Z}_{2}$-grading and a Koszul sign rule. Unless otherwise note, the definitions in this section follow [4].
Definition 3.1.
Let $\mathrm{Ab}_{\mathbb{Z}_{2}}$ be the category of abelian groups with a $\mathbb{Z}_{2}$-grading, which we equip with the symmetric monoidal structure given by the braiding isomorphism
(5) $\begin{split}M\otimes N&\longrightarrow N\otimes M\\ m\otimes n&\longmapsto(-1)^{|m||n|}n\otimes m,\end{split}$
where $|m|$ denotes the degree of the element.
Remark 3.2.
We will be using both $\mathrm{deg}(-)$ and $|\cdot|$ to denote the degree map.
Definition 3.3 ([35]).
A superring $A$ is a $\mathbb{Z}_{2}$-graded ring $A=A_{0}\oplus A_{1}$ such that the product map $A\times A\to A$ satisfies $A_{i}A_{j}\subset A_{i+j}$. A morphism of superrings is a $\mathbb{Z}_{2}$-grading preserving morphism of rings. The elements of $A_{0}$ are called even, the elements of $A_{1}$ are called odd and an element that is either even or odd is said to be homogeneous.
Remark 3.4.
Equivalently, we can define a superring as a commutative monoid object with respect to the symmetric monoidal structure in the category $\mathrm{Ab}_{\mathbb{Z}_{2}}$.
Example 3.5.
Let $k$ be a field and $\{e_{1},...,e_{n}\}$ be a set. We can construct the free odd superalgebra by
(6) $\Lambda_{n}\coloneqq\bigwedge(e_{1},..,e_{n}).$
This algebra is called Graßmann algebra.
Definition 3.6.
A locally ringed superspace is a pair $X=(|X|,\mathcal{O}_{X})$, where $|X|$ is a topological space and $\mathcal{O}_{X}$ is a sheaf of superrings, such that for every point $x\in|X|$ $\mathcal{O}_{X,x}$ is a local superring.
Definition 3.7.
We define the category of affine superschemes to be
(7) $\mathrm{sAff}\coloneqq\mathrm{SRings}^{\mathrm{op}}.$
Definition 3.8.
A superscheme is a locally ringed superspace, which is locally isomorphic to the superspectrum of a superring.
Remark 3.9.
We can interpret the odd coordinates geometrically as fat points. Let $A$ be a superring, then the odd elements $\theta\in A_{1}$ satisfy the relation $\theta^{2}=0$. Let us now consider a superalgebra generated by a single degree 1 element over $\mathbb{C}$, i.e. $\mathbb{C[\theta]}$. We observe that if we forget the $\mathbb{Z}_{2}$-grading, there is an isomorphism
(8) $\mathbb{C[\theta]}\cong\mathbb{C}[\epsilon]/\epsilon^{2}.$
By this analogy, we can think of odd coordinates geometrically as thickening of the point or a first order Taylor expansion around the origin. Moreover, for a general superring $A$ we can also observe that, as topological spaces, we have a homeomorphism
(9) $\mathrm{Spec}(A)\cong\mathrm{Spec}(A/J).$
Definition 3.10.
Let $X$ be a superscheme, then we denote the sheaf generated by all degree 1 nilpotents by $J_{X}$. A superscheme X is called projected if the structural sequence
[…]
splits.
Remark 3.11.
In the category of supermanifolds, it turns out that every real supermanifold is graded in a non-natural way, due to Batchelor’s theorem [2]. So if we want to look for non-projected objects, we must look at complex supermanifolds.
Definition 3.12.
Let X be a superscheme of odd dimension m, then there exists a $J_{X}$-adic filtration of $\mathcal{O}_{X}$
(10) $J_{X}^{0}\supset J_{X}^{1}\supset...\supset J_{X}^{m+1}=0.$
We can now define the sheaf
(11) $\mathrm{Gr}(\mathcal{O}_{X})\coloneqq J_{X}^{0}/J_{X}^{1}\oplus J_{X}^{1}/J_{X}^{2}\oplus...\oplus J_{X}^{m-1}/J_{X}^{m}\oplus J^{m}_{X}$
The graded superscheme associated to X is defined as
(12) $\mathrm{Gr}(X)\coloneqq(|X|,\mathrm{Gr}(\mathcal{O}_{X})).$
Definition 3.13.
A superscheme $X$ is called graded, if there exists an isomorphism $X\cong\mathrm{Gr}(X)$.
Remark 3.14.
The terminology in supergeometry is not yet set in stone. Some authors refer to this property as split [4], [32] and some authors call this property being a graded supermanifold [28]. We remind the reader to always carefully check, what is meant by the term graded or split when encountering supergeometry. It is also an unfortunate fact that all of these words are very overused in a multitude of contexts.
3.2. Algebraic Supergroups
Here we will introduce algebraic supergroups, following [20] and [21], the analogue of algebraic groups in supergeometry. While algebraic supergroups frequently arise in the context of supersymmetric physical models, their fundamental mathematical significance is perhaps best illustrated by Deligne’s theorem on tensor categories.
Theorem 3.15 ([8]).
Every $k$-tensor category $\mathcal{A}$ such that:
  1. (1)
    $k$ is an algebraically closed field of characteristic zero.
  2. (2)
    $\mathcal{A}$ is of subexponential growth.
is a neutral super Tannakian category, and there exists:
  1. (1)
    an affine algebraic supergroup $G$ whose algebra of functions $\mathcal{O}(G)$ is a finitely generated $k$-algebra.
  2. (2)
    a tensor-equivalence of categories
    $\mathcal{A}\simeq\operatorname{Rep}(G,\epsilon)$
    between $\mathcal{A}$ and the category of finite-dimensional representations of $G$.
This theorem roughly states that under mild conditions every tensor category can be realized as the category of representations of some algebraic supergroup.
In the standard literature, schemes are described as locally ringed spaces, whose structure sheaf encodes the algebraic data. Similarly, algebraic groups can be studied as locally ringed space. To make the group axioms compatible with the structure sheaf, we formulate the group axioms as a collection of commutative diagrams. This procedure results in the following
Definition 3.16.
Let $G$ be an $S$-superscheme and consider $m:G\times G\rightarrow G$ to be a morphism of schemes. Furthermore, if there exists morphisms of superschemes
$e:*\rightarrow G\quad,\quad\operatorname{inv}:G\rightarrow G$
such that the following diagrams commute
[…]
then we call $G$ an $S$-group superscheme, where we take the fiber products over $S$. If G is a $k$-superscheme, we call it an algebraic supergroup.
Remark 3.17.
We notice that since an algebraic supergroup is defined by limit diagrams (in this case pullbacks), that whenever we have a limit preserving functor, that group objects will be sent to group objects.
Let $\mathcal{C}$ be an arbitrary category with all small limits, then we call an object $G\in\mathcal{C}$ that satisfies all these conditions a group object. Since the category of superschemes is complete, we obtain, by cheating, an equivalent definition.
Definition 3.18.
An algebraic supergroup $G$ is a group object in the category of superschemes.
Remark 3.19.
Generally, one denotes the group objects inside of a category $\mathcal{C}$ that contains all small limits by $\operatorname{Grp}(\mathcal{C})$.
Another standard way to think about superschemes, is in terms of representable functor, in particular, in this setting one can think of an algebraic group as a functor from
$\mathcal{F}:\operatorname{sSch}^{\operatorname{op}}\rightarrow\operatorname{Grp},$
which we usually call a group functor.
Example 3.20 ([7]).
Let $V$ be a super vector space. Then the functor of points of the associated affine space is
[…]
If we now consider the $R$-points, where $R$ is a superalgebra. Consider a purely odd vector, it is of the form
(14) $v=\sum\alpha_{i}\otimes e_{i}.$
Example 3.21.
Consider $SL_{2}=\operatorname{Spec}(\mathbb{Z}[a,b,c,d]/(ad-bc-1))$. Then its group functor is given by
$\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[a,b,c,d]/(ad-bc-1),-)$
We can now compute
$\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[a,b,c,d]/(ad-bc-1),\mathbb{C})=SL_{2}(\mathbb{C})$
because the morphisms are determined by the evaluation maps for each variable. The evaluations live inside of $\mathbb{C}$ and have to satisfy the determinant condition. The result is precisely $\operatorname{SL}_{2}(\mathbb{C})$.
However, for now let us return to the first definition of a group superscheme. Assume $G$ to be an affine $k$-superscheme and apply $\mathrm{Spec}(-)$ to this collection of diagrams, and we obtain the following:
[…]
So we have a co-group structure on our $k$-superalgebra, which is then called a Hopf superalgebra.
Remark 3.22.
By the functor of points viewpoint a superscheme is already determined by its functor on affines, i.e.
$\operatorname{sRings}\rightarrow\mathrm{Set}.$
If $G$ is an algebraic group over $k$ and $\Lambda$ a $k$-Graßman algebra, then by functoriality we obtain a map
$G(\Lambda)\rightarrow G(k)=G_{red}(k),$
which arises out of the natural quotient map of $\Lambda\rightarrow k$ that sends all odd elements to zero.
Another important notion in the theory of algebraic supergroups, while not immediately relevant here, is the notion of a Harish-Chandra pair. We choose to write this short section as we believe it is quite helpful in general when thinking about algebraic supergroups. It turns out that this category is equivalent to the category algebraic supergroups [11], [20].
Definition 3.23.
Suppose $(G_{0},\mathfrak{g})$ are respectively a group (analytic or affine algebraic) and a super Lie algebra. Assume that:
  1. (1)
    $\mathfrak{g}_{0}\simeq\mathrm{Lie}(G_{0})$,
  2. (2)
    $G_{0}$ acts on $\mathfrak{g}$ and this action restricted to $\mathfrak{g}_{0}$ is the adjoint representation of $G_{0}$ on $\mathrm{Lie}(G_{0})$. Moreover the differential of such action is the Lie bracket. We shall denote such an action with $\mathrm{Ad}$ or as $g.X$, $g\in G_{0}$, $X\in\mathfrak{g}$.
Then $(G_{0},\mathfrak{g})$ is called a super Harish-Chandra pair (SHCP).
Let $\mathrm{SHCP}$ be the category of super Harish-Chandra pairs and $\mathrm{ASP}$ the category of algebraic supergroups then as mentioned earlier we also have the following equivalence of categories
(15) $\mathrm{SHCP}\cong\mathrm{ASP}$
due to Gavarini [11] and proven later again over commutative rings by Masuoka and Shibata in [20].
Remark 3.24.
In [11] the equivalence is essentially constructed by exponentiating all odd elements of the Lie superalgebra. This is well-defined, because due to the nilpotence of odd elements the exponential will truncate at first order. As an example consider $\mathbb{C}[\theta]$ with $\theta^{2}=0$ thus we have.
(16) $\operatorname{exp}(\theta)=1+\theta$
By the same token, we will also see that the Hopf algebra, that corresponds to the algebraic supergroup, can be decomposed as a superscheme into
(17) $k[G]=k[G_{0}]\otimes\bigwedge\mathfrak{g}_{1}^{\vee}$
Remark 3.25.
The group structure is not necessarily preserved under this isomorphism, however the co-unit is preserved [20].
The supergroups whose group structure is preserved under the associated graded functor, were also already classified. It turns out it is equivalent to $\mathfrak{g}_{1},\mathfrak{g}_{1}]=0 [$32].
$3.$ $\operatorname{OSp}(1|2)$
Let $V\coloneqq\mathbb{R}^{1|2}$ be a super vector space over the Grassmann-algebra $\Lambda$ and $B$ be an even non-degenerate bilinear form.
Definition 3.26 ([16]).
Consider the vector space $V$ and $B$ as above, then
(18) $\operatorname{OSp}(1|2)\coloneqq\{g\in\operatorname{GL}(V)|B(gv,gw)=B(v,w)\ \forall v,w\in V\}$
The group law is given by ordinary matrix multiplication. Alternatively, we can also describe it as a group preserving a supersymplectic structure $\omega$ on $\mathbb{R}^{2|1}$ [24]. The descriptions are equivalent and for now we will proceed with the symplectic description. Generally, we will be working over $\mathbb{C}$ and denote this by $\operatorname{OSp(1|2)}$. We can also think of $\operatorname{OSp}(1|2)$ as a group functor from the category of algebraic superschemes to the category of groups. We recover the previous notion of $\operatorname{OSp}(1|2)$-matrices via the $\Lambda$-points of $\mathrm{OSp}(1|2)$, where $\Lambda$ is a $\mathbb{C}$-Graßmann algebra. Explicitly, the $\Lambda$-points are given by matrices of the following form:
$\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}$
with $a,b,c,d,f\in\Lambda_{0}$ and $\alpha,\beta,\gamma,\delta\in\Lambda_{1}$.
Remark 3.27.
This viewpoint is implicitly taken in [25].
These matrices are also subject to some relations, which can be extracted from the symplectic viewpoint for $\mathrm{OSp}(1|2)$. The symplectic form that has to be preserved can be written as a matrix $J$ of the form:
$J=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix}$,
which needs to be invariant under the action of $\operatorname{OSp}(1|2)$. We also obtain a formula for the inverse of an element $g$. Let $g\in\operatorname{OSp}(1|2)$, then its inverse is given by
$g^{-1}=J^{-1}g^{st}J,$
which in formulas is
$g=\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\quad\quad g^{-1}=\begin{pmatrix}d&-b&\delta\\ -c&a&-\gamma\\ -\beta&\alpha&f\end{pmatrix}$.
By explicit calculation, one can also see that these properties force the relations
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc.$
onto us and the relation $\operatorname{Ber}(g)=1$ is imposed separately.
Remark 3.28.
We recall that the functor $(-)_{\mathrm{red}}$ preserves group objects, and one observes that the underlying bosonic group is $\operatorname{SL}_{2}(-)$. The reduction functor induces a morphism of schemes, which is a closed immersion, i.e.
[…]
Let $\iota:k\rightarrow\Lambda$ be the inclusion morphism and $f:\Lambda\rightarrow k$. By functoriality, we obtain the following commutative diagram:
$\mathrm{SL}_{2}(\Lambda)$$\mathrm{OSp}_{1|2}(\Lambda)$$\mathrm{OSp}_{1|2}(k)=\mathrm{SL}_{2}(k)$$\mathrm{SL}_{2}(k)$$\mathrm{OSp_{1|2}(k)}$$\scriptstyle{f_{*}}$$\scriptstyle{\iota_{*}}$$\scriptstyle{\iota_{*}}$
and observe the splitting sequence
[…]
which in formulas is given by
(20) $\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}&0\\ c_{\#}&d_{\#}&0\\ 0&0&1\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}$
We can once again introduce notions like semisimple or reductive, however being a semisimple algebraic supergroup is a very special phenomenom. The classification of semisimple algebraic supergroups is well known and was originally proven by Hochschild and Djoković [9]. Recently Sherman reproved this with more modern techniques in [29]. We will state Sherman’s version.
Theorem 3.29 ([29]).
Let $G$ be a connected algebraic supergroup. Then $Rep(G)$ is semisimple if and only if
(21) $G=K\times\operatorname{OSp}(1|2n_{1})\times...\times\operatorname{OSp}(1|2n_{k})$
for a reductive algebraic group $K$ and positive integers $n_{1},...n_{k}$.
Remark 3.30.
We can see that in the special case we will be tackling, we are still working with a reductive algebraic supergroup, which has nice properties with respect to taking quotients.
Next we describe the mentioned triangulation lemma for $\operatorname{OSp}(1|2)$. Let $\Lambda_{i}$ be the free odd superalgebra in $i$-generators. We will work in $\Lambda$-points, where $\Lambda\coloneqq\underset{i}{\mathrm{colim}}\ \Lambda_{i}$.
Lemma 3.31.
Let $A_{0}\in SL(2,\mathbb{C})\subset\operatorname{OSp}(1|2)$ and $B_{0}\in\operatorname{OSp}(1|2)$. Furthermore, assume that $A$ can be diagonalized as $A=\mathrm{diag}(\mu,\mu^{-1},1)$ and $\mu\neq\mu^{-1}$, then one can always find an element $g\in\operatorname{OSp}(1|2)$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
$gA_{0}g^{-1}=\begin{pmatrix}\mu&0&0\\ 1&\mu^{-1}&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\psi\\ \psi&0&1\end{pmatrix}\quad,\quad gB_{0}g^{-1}=\begin{pmatrix}\lambda&\kappa&0\\ 0&\lambda^{-1}&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&\xi\\ 0&1&0\\ 0&-\xi&1\end{pmatrix}$
Proof.
Let
$g=\begin{pmatrix}0&-x^{-1}&0\\ x&y&\nu\\ 0&-x^{-1}\nu&1\end{pmatrix}$,
where $x,y,\nu$ are indeterminates and conjugate $diag(\mu,\mu^{-1},1)$, which yields
$\begin{pmatrix}\mu^{-1}&0&0\\ xy(\mu-\mu^{-1})&\mu&\nu(-\mu+1)\\ \nu(\mu^{-1}-1)&0&1\end{pmatrix}$
define $\psi\coloneqq\nu(\mu^{-1}-1)$ and choose $xy=(\mu-\mu^{-1})^{-1}$ and we can reduce this to
$A=\begin{pmatrix}\mu^{-1}&0&0\\ 1&\mu&\psi\mu\\ \psi&0&1\end{pmatrix}$.
On the other hand, we have to conjugate the matrix $B_{0}$:
$CB_{0}C^{-1}=\begin{pmatrix}0&-x^{-1}&0\\ x&y&\nu\\ 0&-x^{-1}\nu&1\end{pmatrix}\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\begin{pmatrix}y&-x^{-1}&-x^{-1}\nu\\ x&0&0\\ -\nu&0&1\end{pmatrix}$
The result is rather non enlightening. However, we can use the same reasoning as in the $\operatorname{SL}(2,\mathbb{C})$ case and set the bosonic bottom left corner to be zero. So let us look at the argument and write
$B=CB_{0}C^{-1}=\begin{pmatrix}S&T&\epsilon\\ P&Q&\eta\\ S\eta-P\epsilon&T\eta-Q\epsilon&1+\epsilon\eta\end{pmatrix}$,
where $P=-bx^{2}+cy^{2}+axy-dxy+\alpha\nu+y\beta\nu$. Now set $P=0$, because as previously seen we did not make a choice for $y$ yet, so we can find a solution to that polynomial. This reduces our problem to
$B=\begin{pmatrix}S&T&\epsilon\\ 0&Q&\eta\\ S\eta&T\eta-Q\epsilon&1+\epsilon\eta\end{pmatrix}$,
where we notice, that $\eta$ is a function of $\beta$, so we work with the object $\eta(\nu)$. The function is naturally linear in $\nu$, so we can simply solve for $\nu$ and due to the upper triangular shape we also rewrite $S=\lambda$, $T=\kappa$ and $Q=\lambda^{-1}$ and we arrive at
$B=\begin{pmatrix}\lambda&\kappa&\epsilon\\ 0&\lambda^{-1}&0\\ 0&\lambda^{-1}\epsilon&1\end{pmatrix}$.
Finally, define $\xi\coloneqq\lambda^{-1}\epsilon$ can write down the normal form
$\begin{pmatrix}\lambda&\kappa&\lambda\xi\\ 0&\lambda^{-1}&0\\ 0&\xi&1\end{pmatrix}$.
This explicit computation, now lets us write down ”invariant” odd functions $\theta,\xi$ (invariant up to $\pm$). At this point we have to write down what $\nu$ is. In the conjugation one computes that
$\epsilon$ $=-x^{-1}\beta$
$\eta$ $=-x^{-1}(xa+yc)\beta+(x\alpha+y\nu+\beta f)$
setting $\eta=0$ one can solve for $\nu$ which yields
$\nu$ $=y^{-1}(x^{-1}(xa+yc)\beta-x\alpha-\beta f).$
Collecting all of the previous results we can write down
$\psi$ $=y^{-1}(x^{-1}(xa+yc)\beta-x\alpha-\beta f)(\mu^{-1}-1)$
$\xi$ $=-\lambda^{-1}x^{-1}\beta,$
where we remind ourselves that $x=y^{-1}(\lambda-\lambda^{-1})^{-1}$ and $y$ is a solution of the polynomial $P$. We are working in a non-integral ring, so the existence of a solution is non-obvious. However, a result of Shigeaki Nagamichi [14], tells us as long as the reduction possesses roots, the corresponding super polynomial also has roots. ∎
Remark 3.32.
Although it is proven for complex coefficients, we can see that the same proof strategy works for the case of $\operatorname{OSp}(1|2)$ with real coefficients, where the reduction of that element is hyperbolic.
Remark 3.33.
In [31] they make use of this parametrization of the moduli space of flat $\operatorname{OSp}(1|2)$-connections to compute an explicit form of $G$-invariant measures. They also write down the pair in the form
$U$ $=\begin{pmatrix}e^{a}&\kappa&0\\ 0&e^{-a}&0\\ 0&0&1\end{pmatrix}exp(\xi q_{1})$ $V=\begin{pmatrix}e^{-b}&0&0\\ 1&e^{b}&0\\ 0&0&1\end{pmatrix}exp(\xi q_{2})$
$q_{1}$ $=\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&-1&0\end{pmatrix}$ $q_{2}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$.
This might look different compared to the parametrization given earlier, but one recalls that the Taylor expansion of odd elements vanishes after the first order. Expanding it that way gives you the desired result. Furthermore, one notes that this parametrization of our pair $(A,B)$ ensures that there are 3 even degrees of freedom and 2 odd degrees of freedom (defined up to sign). We also note that $q_{1},\ q_{2}$ are the generators of $\mathfrak{g}_{1}$ of $\mathfrak{osp}(1|2)$ and by exponentiation they become elements in the corresponding algebraic supergroup.
3.4. Invariant Theory
In this section, we begin by reproving the Fricke-Klein case again with a method that directly aligns with the method of our main result. This explicit worked example should help elucidate the proof strategy for the supercase. Afterwards, we introduce the necessary tools from classical invariant theory and adapt them to the $\mathbb{Z}_{2}$-graded case. Most of these results will generalize without adapting the proof and can be skipped for the experienced reader. We follow [15] and [23].
First a short remark on $\operatorname{OSp}(1|2)$. Since we have semisimple representations for $\operatorname{OSp}(1|2)$ we also get that it is linearly reductive. We start by introducing the notion of a Reynolds operator.
Definition 3.34 ([23]).
Let $R:V\rightarrow V$ be a linear map of a vector space $V$. Let $G$ be an algebraic group G, which acts on V rationally. It is called a Reynolds operator if $R$ is a projection onto $V^{G}$, the space of invariant vectors in $V$.
Suppose this operator exists and let $Y$ be a subvariety of $X$. Then we have a short exact sequence of $G$-modules, which by linear reductivity gives us also a splitting and hence the diagram
[…]
The induced map on invariant subrings is surjective by the Reynolds property and we can check injectivity by diagram chasing.Therefore,
(22) $k[Y]^{G}=k[X]^{G}/I^{G}.$
Let us employ the outline strategy in a specific example, i.e. let us reprove the theorem of Fricke-Klein again.
Example 3.35.
Consider the ring $\mathbb{C}[M_{2}^{2}]^{SL_{2}}$, the ring of invariants of $2\times 2$ matrices invariant under conjugation by $SL_{2}$. It is known that this ring is
(23) $\mathbb{C}[M_{2}^{2}]^{SL_{2}}=\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB),\operatorname{tr}(A^{2}),\operatorname{tr}(B^{2})].$
the ideal $I$ is generated in this case by $I=\langle\operatorname{det}(A)-1,\operatorname{det}(B)-1\rangle$. At first one might think, that this will not quotient out any generators but we can remedy that by a simple application of the Cayley-Hamilton theorem. For a $2\times 2$ matrix $A$ it gives us the equation:
(24) $A^{2}-\operatorname{tr}(A)A+\operatorname{det}(A)\operatorname{id}=0,$
which after applying the trace to the whole equation gives us
(25) $\begin{split}\operatorname{tr}(A^{2})-\operatorname{tr}(A)^{2}+2=0\\ \operatorname{tr}(A^{2})=\operatorname{tr}(A)^{2}-2.\end{split}$
This now makes the generators $\operatorname{tr}(A^{2}),\operatorname{tr}(B^{2})$ redundant and we get
(26) $\mathbb{C}[M_{2}^{2}]^{SL_{2}}/I^{SL_{2}}=\mathbb{C}[\operatorname{tr}(A),\operatorname{tr}(B),\operatorname{tr}(AB)].$
One should be careful though, that this isomorphism is an isomorphism of modules and not of rings. However, it does give us enough information about the number of generators which suffices.
Remark 3.36.
Even though this proof used the Cayley-Hamilton theorem, it is different to the proof given in [12].
Let us establish, that indeed for a linearly reductive supergroup, we can once again define the Reynolds operator.
Theorem 3.37.
Let $G$ be a linearly reductive algebraic supergroup that acts rationally on a $k$-vector space $V$ and let $V^{G}$ be the subspace of invariant vectors. Then there exists a uniquely defined linear operator
(27) $R:V\longrightarrow V$
projecting $V$ onto $V^{G}$.
Proof.
If $v\in V$ is not invariant, then it is contained in some subspace $W$, such that $G\cdot W\subset W$. By linear reductivity $W$ once again splits into
(28) $W=W^{G}\oplus W^{\prime}$
It follows that $v\in W^{\prime}$, so it is non-empty invariant and $W^{\prime}\cap V^{G}=0$. By applying Zorn’s lemma to the set of invariant subspaces, with this property we can find a maximal one, which we denote by $V^{\prime}$. Suppose that there are two such maximal subspaces denoted by $V_{1},V_{2}$. It holds for the direct sum that $V_{1}\oplus V_{2}\cap V^{G}=0$. To show that, assume that it is non-empty, thus we can find an element $w=v_{1}\oplus v_{2}$, which is invariant under the group action. However, since the action acts on both parts separately, this is a contradiction, hence $V_{1}\oplus V_{2}\cap V^{G}=0$. This means, that the direct sum, would be a larger subspace with the wanted property, which once again is a contradiction. Thus $V^{\prime}$ is uniquely determined and we can also see that
(29) $V^{G}\oplus V^{\prime}=V.$
Now we define $R$ by the property, that it fixes $V^{G}$ and has $V^{\prime}$ as its kernel. ∎
Remark 3.38.
From here on out, we could prove Nagata’s theorem in the supercase, however that is not really useful, since previously we have seen that the class of redudctive supergroups is not very large. Therefore, it is not nearly as powerful as in the case of algebraic groups.
Remark 3.39.
One can prove, that the existence of a Reynolds operator is equivalent to the representation category being semisimple. It is also equivalent to the fixed point functor/invariants functor $(-)^{G}$ being exact.
We continue by establishing that polarization and restitution works just as in the purely bosonic case. For the reader unfamiliar with classical invariant theory, polarization allows us to transform the polynomial into a multilinear expression, where we already know the invariants and restitution is the inverse process. We combine in the following way
PolarizationComputing InvariantsRestitution
which allows us to compute the polynomial invariants, which arise out of the multilinear ones. To be a little bit more precise, we want to show that
$\bigoplus_{d}k[V]_{d}$$\bigoplus_{d}k[V^{d}]$$\bigoplus k[V]_{d}$$\scriptstyle{\mathcal{P}}$$\scriptstyle{F}$
is $G$-equivariant, where the maps will be explained shortly. We can see that one can simply chase an invariant polynomial $f\in k[V]$ through this construction, find an explicit form in terms of multilinear expressions and then recover the original polynomial in terms of the multilinear invariants. This strategy shows, why it is enough to be able to compute multilinear invariants, which is much easier compared to the more difficult problem of finding polynomial invariants right off the bat.
Definition 3.40.
Consider a direct sum $V=V_{1}\oplus...\oplus V_{r}$ of finite dimensional super vector spaces. A function $f\in k[V_{1}\oplus...\oplus V_{r}]$ is called mltihomogeneous of degree $h=(h_{1},...,h_{r})$ if $f$ is homogeneous of degree $h_{i}$ in $V_{i}$, i.e. for all $v_{1},...,v_{r}\in V$ and $t_{1},...,t_{r}\in k$ we have
(30) $f(t_{1}v_{1},...,t_{r}v_{r})=t_{1}^{h_{1}}\cdot...\cdot t_{r}^{h_{r}}f(v_{1},...v_{r}).$
One can also call $h$ the multidegree of $f$.
Every polynomial function $f$ is in a unique way a sum of multihomogeneous functions $f=\sum f_{h}$. The $f_{h}$ are usually called multihomogeneous components of $f$. This gives rise to a decomposition
(31) $k[V_{1},...,V_{r}]=\bigoplus_{h\in\mathbb{N}^{r}}k[V_{1},...,V_{r}]_{h},$
where $k[V_{1},...,V_{r}]_{h}$ denotes the subspace of multihomogeneous functions of degree $h$.
Remark 3.41.
This is a different grading from the natural $\mathbb{Z}_{2}$-grading of our superring.
Remark 3.42.
The way this interacts with super is demonstrated in the following. Assume $V_{r}$ is entirely odd, then it means the polynomial $f$ can be at most homogeneous of degree $h_{r}=1$
It is clear that the multidegree is preserved under the action of $\operatorname{GL}(V_{1})\times...\times\operatorname{GL}(V_{r})$. Furthermore, if $V_{1},...V_{r}$ are representations, then we also have a natural grading on the ring of invariants
(32) $k[V_{1},...,V_{r}]^{G}=\bigoplus_{h\in\mathbb{N}^{r}}k[V_{1},...,V_{r}]_{h}^{G}.$
Definition 3.43.
Let $V$ be a finite dimensional $k$-vector space. Identify $k[V]=\bigoplus_{d}k[V]_{d}$, where $k[V]_{d}$ denotes the space of homogeneous polynomials of degree $d$. Let $f\in k[V]_{d}$ and $v=\sum_{i}^{d}t_{i}v_{i}$ for $v_{i}\in V$. We define
$\mathcal{P}(f)(\sum_{i}t_{i}v_{i})=\sum_{s_{1}+...+s_{d}=d}t_{1}^{s_{1}}...t_{d}^{s_{d}}f_{s_{1},...,s_{d}}(v_{1},...,v_{d}),$
where $f_{s_{1},...,s_{d}}\in k[V^{d}]$ are well-defined and multihomogeneous of multidegree $(s_{1},...,s_{d})$. We call $f_{1,...,1}\in k[V^{d}]$ the (full) polarization of $f$ denoted by $\mathcal{P}f$.
Lemma 3.44.
The linear operator $\mathcal{P}\colon K[V]_{d}\to K[V^{d}]_{(1,1,\ldots,1)}$ has the following properties:
  1. (a)
    $\mathcal{P}$ is $\mathrm{GL}(V)$-equivariant;
  2. (b)
    $\mathcal{P}f$ is symmetric;
  3. (c)
    $\mathcal{P}f(v,v,\ldots,v)=d!\,f(v)$.
Proof.
For equivariance consider the diagram
$k[V]_{d}$$k[V^{d}]$$k[V]_{d}$$k[V^{d}]$$\scriptstyle{\mathcal{P}}$$\scriptstyle{g\cdot}$$\scriptstyle{g\cdot}$$\scriptstyle{\mathcal{P}}$
and let
(33) $f(t_{1}v_{1}+...+t_{d}v_{d})=\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(v_{1},...v_{d}).$
We start chasing
$\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(v_{1},...v_{d})$$t_{1}\cdot...\cdot t_{d}f_{s_{1}...s_{d}}(v_{1},...v_{d})$$\sum t_{1}^{s_{1}}\cdot...\cdot t_{d}^{s_{d}}f_{s_{1}...s_{d}}(g^{-1}v_{1},...,g^{-1}v_{d})$$t_{d}f_{s_{1}...s_{d}}(g^{-1}v_{1},...g^{-1}v_{d})$$\scriptstyle{g\cdot}$$\scriptstyle{g\cdot}$
$\mathcal{P}$ is also obviously symmetric. It remains to show $(c)$. Let us calculate
(34) $f(t_{1}v,...,t_{d}v)=(\sum_{i}t_{i})^{d}f(v)=(d!t_{1}\cdot...\cdot t_{d}+...)f(v).$
Looking at the first factor the claim follows. ∎
Definition 3.45.
For a multilinear $F\in K[V^{d}]$ the homogeneous polynomial $\mathcal{R}F$ defined by $\mathcal{R}F(v):=F(v,v,\ldots,v)$ is called the (full) restitution of $F$.
By construction we also have the following result.
Proposition 3.46.
Assume $\operatorname{char}k=0$ and let $V$ be a finite dimensional representation of a supergroup $G$. Then every homogeneous invariant $f\in K[V]^{G}$ of degree $d$ is the full restitution of a multilinear invariant $F\in K[V^{d}]^{G}$.
Proof.
Again, $\mathcal{R}\colon K[V^{d}]_{(1,\ldots,1)}\to K[V]_{d}$ is a linear $\mathrm{GL}(V)$-equivariant operator and we have $\mathcal{R}\mathcal{P}f=d!f$ by the property (c) of the lemma above. In fact, $f$ is the full restitution of $\frac{1}{d!}\mathcal{P}f$, which is a multihomogeneous invariant. ∎
We continue building up all necessary results, now we will begin describing how to translate vector invariants to matrix invariants. Firstly, we will state the first fundamental theorem for $GL(V)$
Theorem 3.47.
Let $V$ be a finite dimensional super vector space. The multilinear invariants under the natural $GL(V)$-action are given by
(35) $\begin{split}V^{\otimes N}\otimes(V^{\vee})^{\otimes N}&\longrightarrow\mathbb{C}\\ v_{1}\otimes...\otimes v_{N}\otimes\phi_{1}\otimes...\otimes\phi_{N}&\longmapsto\prod_{i}\langle v_{i},\phi_{\sigma(i)}\rangle\end{split}$
and all its transforms under the symmetric group
It remains to show, that the pairings are still given by traces, just like in [26].
Proposition 3.48.
Let V be a $p|q$ dimensional super vector space and $V^{\vee}$ its dual. We take $\langle\cdot,\cdot\rangle:V\otimes V^{\vee}\rightarrow k$, then
$\langle v,\phi\rangle=\operatorname{str}(v\otimes\phi)$
Proof.
We prove this by calculation. Take a homogeneous basis $\left\{e_{j}\right\}$ for $V$ and let $\left\{\varphi_{j}\right\}$ be its dual basis. We take a vector $v\otimes\varphi\in V\otimes V^{\vee}$ and compute the coefficients of the associated matrix $A$. Let
$v$ $=\sum_{i}a_{i}e_{i}\quad,\ \text{such that}\ \forall|e_{i}|=\{0,1\}:\ a_{i}\in\Lambda_{\{0,1\}}$
$\varphi$ $=\sum_{j}b_{j}\varphi_{j}\quad,\ \text{such that}\ \forall|\varphi_{j}|=\{0,1\}:\ b_{j}\in\Lambda_{\{0,1\}}$
and we compute
$v\cdot\varphi=\sum_{ij}a_{i}b_{j}e_{i}\otimes\varphi_{j}$
on the other hand we can also compute
$\langle v,\varphi\rangle=\varphi\cdot v=\sum_{i}(-1)^{|e_{i}||\varphi_{i}|}a_{i}b_{i}.$
It is easy to see that
$\operatorname{str}(A)=\langle v,\varphi\rangle$
Lemma 3.49.
Let $V$ be a super vector space. Then the space of multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is isomorphic to the space of multilinear functions on $\operatorname{End}(V)^{m}$
Proof.
The space of multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is given by
(36) $(V^{\otimes m}\otimes(V^{\vee})^{\otimes m})^{\vee}\cong((V\otimes V^{\vee})^{\otimes m})^{\vee}\cong(\operatorname{End}(V)^{\otimes m})^{\vee}.$
We immediately recognize that the right hand side are the multilinear functions on $\operatorname{End}(V)^{m}$. ∎
Remark 3.50.
All these spaces naturally admit a $\operatorname{GL}(V)$-action. Considering all these vector spaces as $\operatorname{GL}(V)$-representations, it is also easy to see, that these identifications are a $GL(V)$-equivariant isomorphism.
We can now translate the multilinear invariant in vector form into the multilinear invariant in matrix form. Consider $\operatorname{End}(V)^{\otimes i}$ and choose $\sigma\in Sym_{i}$. Decompose $\sigma$ into disjoint cycles
(37) $\sigma=(i_{1}i_{2}...i_{k})(j_{1}j_{2}...j_{n})...(t_{1}...t_{e})$
including the ones of length 1.
Theorem 3.51.
Given $A_{1},...,A_{i}\in\operatorname{End}(V)$, we have:
(38) $\mu_{\sigma}(A_{1}\otimes...\otimes A_{i})=\operatorname{str}(A_{i_{1}}...A_{i_{k}})\operatorname{str}(A_{j_{1}}...A_{j_{n}})\operatorname{str}(A_{t_{1}}...A_{t_{e}})$
Proof.
Since both sides are multilinear, it is enough to prove it for the case where $A_{i}=\phi_{i}\otimes v_{i}$.
(39) $\mu_{\sigma}(A_{1}\otimes...\otimes A_{i})$ $=\prod_{i}\langle\phi_{\sigma(i)},v_{i}\rangle$
(40) $=\langle\phi_{i_{2}},v_{i_{1}}\rangle\langle\phi_{i_{3}},v_{i_{2}}\rangle...\langle\phi_{t_{1}},v_{t_{e}}\rangle$
Consider, for instance, the product
(41) $M$ $=\langle\varphi_{i_{2}},v_{i_{1}}\rangle\langle\varphi_{i_{3}},v_{i_{2}}\rangle\cdots\langle\varphi_{i_{k}},v_{i_{k-1}}\rangle\langle\varphi_{i_{1}},v_{i_{k}}\rangle.$
One verifies immediately that
(42) $\varphi_{i_{1}}\otimes v_{i_{1}}\cdot\varphi_{i_{2}}\otimes v_{i_{2}},\ldots,\varphi_{i_{k}}\otimes v_{i_{k}}$ $=\langle\varphi_{i_{2}},v_{i_{1}}\rangle\langle\varphi_{i_{3}},v_{i_{2}}\rangle\cdots\langle\varphi_{i_{k}},v_{i_{k-1}}\rangle\varphi_{i_{1}}\otimes v_{i_{k}}.$
Therefore,
(43) $M=\mathrm{tr}(A_{i_{1}}A_{i_{2}}\cdots A_{i_{k}}),$
The last bit of prep work we have to do now is simply state the first fundamental theorem of invariants for $\operatorname{OSp}(1|2)$. Let $V$ be a finite dimensional super vector space with $dim(V)=m|2n$ and let $B$ be the symmetric pairing preserved under $\operatorname{OSp}(V)$. Deligne, Lehrer and Zhang proved a first fundamental theorem for $\operatorname{OSp}(V)$ [7].
Theorem 3.52 ([7]).
Let $V$ be as above, then the multilinear $\operatorname{OSp}(V)$-invariant functions on $V^{N}$ for $N$ even are given by
(44) $\begin{split}V^{\otimes N}&\longrightarrow\mathbb{C}\\ v_{1}\otimes...\otimes v_{N}&\longmapsto B(v_{1},v_{2})\cdot...\cdot B(v_{N-1},v_{N}),\end{split}$
and all its permutations in $Sym_{N}$.
This was also proven by Lehrer and Zhang in [16] with a different method of proof earlier than the stated result mostly in the language of Brauer diagrams. With that method they also computed the polynomial invariants ones from the multilinear ones, which given by the monomials in the multilinear invariants.
Theorem 3.53.
Let $V$ be as above, then the polynomial $\operatorname{OSp}(V)$-invariants are polynomials in the variables $B(v_{i},v_{j})\coloneqq(i,j)$ $\forall 1\leq i<j\leq m$
We can now prove a second fundamental theorem for the polynomial invariants. to do that, we need to prove a small lemma first.
Lemma 3.54.
Let $G\coloneqq(B(v_{i},v_{j}))_{i,j}$ then $\operatorname{det}(G)\neq 0$ iff $v_{1},...v_{N}$ are linearly independent.
Proof.
We proof this by contrapositive. If $\operatorname{det}(G)=0$ then there exists a relation between the rows $R_{i}\coloneqq(B(v_{i},v_{1}),...,B(v_{i},v_{N}))$, i.e.
(45) $\sum_{i}a_{i}R_{i}=0$
for some non-zero $a_{i}$. It follows
(46) $\sum_{i}a_{i}R_{i}=(B(\Sigma_{i}v_{i},v_{1}),...,B(\Sigma_{i}v_{i},v_{N}))=0$
So each factor must be zero, thus consider the linear combination
(47) $\begin{split}a_{1}B(\Sigma_{i}v_{i},v_{1})+...+a_{N}B(\Sigma_{i}v_{i},v_{N})&=0\\ B(\Sigma_{i}v_{i},\Sigma_{i}v_{i})&=0\end{split}$
by non-degeneracy, they are linearly dependent. For the other direction, we also consider the contrapositive. if $v_{1},...,v_{N}$ are linearly dependent. Let $a_{i}$ be the coefficient such that the linear combination of the vectors vanishes, then consider the sum of the rows
(48) $\begin{split}&=\sum_{i}a_{i}R_{i}\\ &=(B(\Sigma a_{i}v_{i},v_{1}),...,B(\Sigma a_{i}v_{i},v_{N}))\\ &=B(B(0,v_{1}),...,B(0,v_{N}))=0\end{split}$
so there is a relation between the rows, thus the determinant must vanish. ∎
Theorem 3.55.
Let everything be as above, then the polynomial ring of $\operatorname{OSp}(V)$-invariants is
(49) $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}/I_{n}$
Proof.
We will be investigating the kernel of the morphism of rings
(50) $\phi:\mathbb{C}[(i,j)]\longrightarrow\mathbb{C}[V^{m}]^{\operatorname{OSp(V)}}.$
Notice, that we can fit all the inner products into a matrix $G\coloneqq(B(v_{i},v_{j}))_{i,j}$. One also sees that $G\in\operatorname{Mat}_{N\times N}(\Lambda)$. By the earlier lemma we follow that the maximal rank of this matrix is $n+2m$. It follows, that all $(n+2m+1)\times(n+2m+1)$ minors must vanish, thus the morphism factors in the following way
$\mathbb{C}[(i,j)]$$\mathbb{C}[V^{m}]^{\operatorname{OSp(V)}}$$\mathbb{C}[(i,j)]/I_{n}$}
It is easy, to see that $I_{n}\subset\operatorname{ker}(\phi)$. Furthermore, let $f\in\operatorname{ker}(\phi)$. We use the previous factoring
[…]
where $I_{n}$ is the ideal generated by all $(n+1)\times(n+1)$ minors. we need to check that $\operatorname{ker}(\phi)\subset I_{n}$, which means we have to determine
(51) $f\operatorname{mod}I_{n}\longrightarrow 0.$
We know that $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}$ is also generated by all $(i,j)$, thus the morphism is injective, hence $f\operatorname{mod}I_{n}=0$, therefore $\operatorname{ker}(\phi)=I_{n}$
(52) $\mathbb{C}[V^{m}]^{\operatorname{OSp}(V)}=\mathbb{C}[(i,j)]/I_{n}$
Now as a corollary we have
Corollary 3.56.
The only relation for the invariant ring of $\mathbb{C}[V^{4}]^{\operatorname{OSp}(1|2)}$ is $\operatorname{det}(G)=0$
Proof.
for $\operatorname{OSp}(1|2)$ and for copies of $V$ that all $4\times 4$ minors must vanish, which means the determinant vanishing is the only relation. ∎
A consequence of this, is that we have 9 independent generators.
4. Main Result
We can now state our main result. Let $V$ be a $1|2$-dimensional super vector space.
Theorem 4.1.
The $\operatorname{OSp}(V)$-invariant polynomial functions (by conjugation) on $\operatorname{OSp}(V)\times\operatorname{OSp}(V)$ has 7 independent generators.
Proof.
Let us consider the algebra
(53) $\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}.$
By polarization, we can carry the polynomial invariants to the multilinear invariants. Now we remember that
(54) $((\operatorname{End}(V)\otimes\operatorname{End})^{\vee})^{\operatorname{OSp}(V)}\cong((V\otimes V\otimes V\otimes V)^{\vee})^{\operatorname{OSp}(V)}.$
By (3.52) we know that the number of polynomial invariants of $V^{\otimes 4}$ is 10. By (3.55) we know that there is one relation for $\operatorname{dim}(V)=1|2$ and thus 9 independent generators. Because the multilinear invariants of matrices and vectors are the same, the polynomial invariants are also the same. By reductivity we have
[…]
where $I$ is the defining ideal of $\operatorname{OSp}(V)^{2}$ as subvariety of $\operatorname{End}(V)^{2}$. It is generated by the relations
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc,$
together with the condition $\operatorname{Ber}(A)=1$. We know that there exist no odd invariants on $\mathbb{C}[\operatorname{End}(V)^{2}]$ and we also know the co-action on $I$ sends the odd generators to odd generators. Thus all the odd relations, do not give us an invariant of $I$. However, the relation $\operatorname{Ber}(A)=1$ is invariant under the co-action, thus it is an invariant vector in $I$. It follows that
(55) $I^{\operatorname{OSp}(1|2)}=\langle\operatorname{Ber}(A),\operatorname{Ber}(B)\rangle.$
We denote by $\mathrm{gen}(-)$ the number of generators of a ring. Then
(56) $\begin{split}\mathrm{gen}(\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)})&=\mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})+\mathrm{gen}(I^{\operatorname{OSp}(V)})\\ 9&=\mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})+2\\ \mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})&=7\end{split}$
By an (3.48), we know that all multilinear invariants are given by traces, thus all these invariants are traces as well. ∎
5. Geometric Interpretations
5.1. Character Varieties
We follow [30]. Let $G$ be a reductive algebraic group over k. When looking at the set $\mathrm{Hom}(\Gamma,G)$, we can associate to it a ring $\mathcal{R}(\Gamma,G)$, and the $k$-points of $\operatorname{Spec}(\mathcal{R}(\Gamma,G))|=\operatorname{Hom}(\Gamma,G)$. Since every finitely generated group can be constructed by imposing a set of relations on the free group $F_{n}$, we will start by constructing the representation variety of $F_{n}$.
Proposition 5.1.
Let $G$ be an algebraic group over $k$ and $F_{n}$ the free group on $n$ letters then coordinate ring of the representation variety is $\mathcal{R}(F_{n},G)=k[G]^{\otimes n}$.
Proof.
It is clear, that a morphism of groups between $F_{n}$ and $G$ is determined by the images of the generators, there are no relations and thus a representation of $F_{n}$ is simply a tuple of $n$ elements of $G$, hence a point in the representation variety is given by
(57) $\mathcal{R}(F_{n},G)=k[G]^{\otimes n}$
If we now impose relations on the free group, we get
$\operatorname{Hom}(\Gamma,G)=\operatorname{Spec}(k[G\times...\times G]/I_{\Gamma}).$
The ideal $I_{\Gamma}$ is generated by the relations induced by $\Gamma$, which tell you that the images of the generators must satisfy the relations of $\Gamma$, because otherwise it would not be a group homomorphism. By the quotient nature of $R(\Gamma,G)$ it is not obvious if it is an integral domain, which over $\mathbb{C}$ allows you to identify Specm and Spec. This space algebraically classifies all so-called $G$-representations. However we do want to quotient out all conjugacy equivalent representations. Thus we define
Definition 5.2.
Let $G$ be a reductive algebraic group over $k$ and $\Gamma$ a finitely generated group, then we call the categorical quotient
(58) $\operatorname{Spec}(R(\Gamma,G)^{G})$
the character variety.
Remark 5.3.
If one considers reductive algebraic groups then it turns out that $\operatorname{Specm}(R(\Gamma,G)^{G})=\operatorname{Hom}(\Gamma,G)$ is a good categorical quotient and by Nagata’s theorem it is finitely generated. One should remark, that non-reduced points may still exist. Let $\operatorname{Hom}^{i}(\Gamma,G)$ be the subset of irreducible representations then
$\operatorname{Hom}^{i}(\Gamma,G)//G$
is a good geometric quotient, i.e. the set corresponds to the set theoretic quotient, where if we now consider fibers of the quotient map, then we will recover the orbits of $G$. In general, there should be no expectation, that the full character variety is actually a good geometric quotient, i.e. it is not the orbit space of the representation variety.
We can easily see, that there are no obstructions to replacing the algebraic group with an algebraic supergroup.
Remark 5.4.
Seeing how this construction actually only cares about the Hopf algebras, we can also generalize this question to quantum groups. This will definitely work if we only consider group relations, that can be manipulated in a way such that no inverses appear, because if we consider the quantum group via algebra-valued matrices, then the inverse does not necessarily live in the same algebra. A construction was already given in the physics literature [1]. We note that for quantum groups one reasons purely in the dual picture of the Hopf algebra, so one has to think about coinvariants of the conjugation coaction.
Using this viewpoint, we can reinterpret 4.1 as the fact that the character variety of the free group on 2 letters of $\operatorname{OSp}(V)$ has 7 independent generators, representing a deviation of the classical result for $SL(2,\mathbb{C})$.
There have been results [25], where in the real analytic case one can prove that this space is actually of the expected real dimension $6|4$. One can easily see, why this fails in the algebraic case. We have seen that all invariants on the space of matrices, are given by a supertrace. We recall that
(59) $\operatorname{str}:\operatorname{OSp}(V)\longrightarrow\mathbb{C}^{1|0}$
and thus we cannot determine any odd generators with the help of traces. It is also a consequence of the fact that there is a natural $\mathbb{Z}_{2}$-action acting on the odd coordinates. Explicitly, it is
$=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&\xi\\ 0&1&0\\ 0&-\xi&1\end{pmatrix}\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$
$=\begin{pmatrix}1&0&-\xi\\ 0&1&0\\ 0&\xi&1\end{pmatrix}$.
So this tells us, that if we take the quotient by the full group $\operatorname{OSp}(1|2)$, that the space has to be of purely bosonic dimension. One also sees by this calculation that taking the full ring of invariants can only give you a non-reduced ordinary scheme, since the odd degrees of freedom are not invariant, as we always have to pair two odd coordinates that will be flipped simultaneously to become an invariant.
This means what is $\operatorname{OSp}(1|2)$-invariant is the product of both odd guys. From this perspective, one should consider the quotient $\operatorname{{POSp}}(1|2)\coloneqq\operatorname{OSp}(1|2)/\mathbb{Z}_{2}$ and consequently, one also needs to consider the character variety of $\mathrm{POSp}(1|2)$. The character variety of $\operatorname{OSp}_{\mathbb{R}}(1|2)$ has already been investigated from the analytic point of view in [25], where Penner-coordinates were generalized. It has also appeared in [31], where they gave an analogous recursion formula to Mirzakhani [22] for super hyperbolic surfaces with heuristics coming from supergeometry. The statement was proven with algebro-geometric methods in [24].
Remark 5.5.
One might be disappointed by the structure of this character variety. However, as it turns out there are also other instances of moduli spaces of super objects that are of purely even nature like the Picard scheme of a SUSY curve [4] or the coarse moduli of a superstack [6]. This suggests that trying to obtain a coarse moduli in supergeometry is ill-advised and one should work purely stack theoretically so you do not accidentally kill all odd directions.
5.2. Character Stack
In view of (4.1) it is clear that to get a proper handle on the character variety of algebraic supergroups one should really be looking at the character stack of the algebraic supergroup instead. We give a brief account on a gradedness result of the character stack of $\mathrm{OSp}(1|2)$. Let $X$ be a super Riemann surface and consider the representation variety
(60) $\mathrm{Rep}(X)\coloneqq\operatorname{Hom}(\pi_{1}(X),\operatorname{OSp}(1|2))$
and its stack quotient $\mathfrak{X}\coloneqq[\mathrm{Rep}(X)/\operatorname{OSp}(1|2)]$, called the character stack. For brevity, we will only define the notion of prestacks and argue, why for this papers purpose it is enough and we will give a short user’s guide on how one works with stacks in practice and for more details we refer to [5].
Definition 5.6.
A pre-superstack is an element in the category
(61) $\mathrm{PresStk}\coloneqq\mathrm{Fun}(\mathrm{sRings},\mathrm{Gpd}),$
i.e. it is a $(2,1)$-presheaf of groupoids.
Remark 5.7.
To turn a pre-superstack into a superstack, one needs to equip it with a Grothendieck topology and it needs to satisfy the descent condition for the chosen topology (this means the cover needs to satisfy a higher categorical gluing condition, in this case a 2-categorical gluing condition).
Conceptually, the thory of stacks is best approached trhough the functor of points perspective. From this perspective the theory of stacks does not look much different from the theory of schemes. We will state two facts that allow us to take a shortcut, so we do not have to think about actual stacks. Firstly, the category of quasi-coherent sheaves on the prestack $\mathfrak{X}$ is equivalent to the category of quasi-coherent sheaves on its stackification $\mathfrak{X}^{+}$
(62) $\mathrm{QCoh}(\mathfrak{X})\cong\mathrm{QCoh}(\mathfrak{X}^{+}).$
Furthermore, the category of superstacks forms a full subcategory of the category of pre-superstacks, so it admits a fully faithful embedding, that is essentially surjective
(63) $\mathrm{sStk}\subset\mathrm{PresStk}.$
This tells us that we do not need to care about the topologies of the stack, if you want to write down morphisms of stacks, the only thing one needs to do is to write down a natural transformation between the functor of points of the stacks and in a similar vein, you do not need to think about the topology to be able to think about quasi-coherent sheaves. Moreover, by its classical definition a stack is also a category fibered in groupoids, so in particularly it is a category, so one only needs to write down a functor to construct a morphism of stacks that commutes with projections to the base category. With all this out of the way, we turn to our short analysis of the character stack.
We want to investigate whether this space is graded, since it is one of the most basic properties one can check. The question whethera space is graded or not depends on a classical obstruction theory argument, i.e. does a certain cohomology class vanish on the underlying space [34] [18]. Thus in this specific setting, it is enough to think about the prestack/functor of points of the character stack. We recall that for a stack quotient $[X/G]$ of a scheme, we have an equivalence of categories
(64) $\operatorname{QCoh}([X/G])\cong\operatorname{QCoh}^{G}(X).$
Furthermore, for $X\in\operatorname{CRing}^{\operatorname{op}}$ and $\mathcal{F}\in\operatorname{QCoh}(X)$, $H^{i}(X,\mathcal{F})=0\ \forall i>0$. We can also rephrase this as the global section functor being exact.
Proposition 5.8.
Let $k$ be an algebraically closed field of characteristic 0, $G$ a reductive algebraic supergroup, and $X$ an affine $k$-superscheme. Then
(65) $H^{i}_{G}(X,\mathcal{F})=0\quad\forall i>0.$
Remark 5.9.
This property is called being cohomologically affine.
Proof.
The global sections functor $\Gamma$ is exact, because $X$ is affine and $(-)^{G}$ is exact, because $G$ is reductive. $G$-equivariant sheaf cohomology is defined to be the right derived functor of the composition $(-)^{G}\circ\Gamma$, thus by composition of exact functors, the composition is also exact. We obtain
(66) $H^{i}_{G}(X,\mathcal{F})\subset H^{i}(X,\mathcal{F})=0\ \forall i>0.$
Corollary 5.10.
Let $X\in\operatorname{Sch}$ and $G\in\operatorname{Grp(Sch)}$. If $X$ is cohomologically affine and smooth and the fixed point functor $(-)^{G}$ is exact, then the stack quotient $[X/G]$ is graded.
Proof.
Let $\mathcal{J}$ be the sheaf of nilpotents on $[X/G]$ and $\mathcal{E}\coloneqq\mathcal{J}/\mathcal{J}^{2}$. Furthermore, we denote $\mathcal{F}^{i}\coloneqq T_{(-)^{i}}[X/G]_{\operatorname{red}}\otimes\wedge^{i}\mathcal{E}^{\vee}$. To prove $[X/G]$ is graded, we need to show the that obstruction classes
(67) $\omega_{i}\in H^{1}([X/G]_{\operatorname{red}},\mathcal{F}^{i})$
vanish [34]. Given the canonical inclusion $\iota:[X/G]_{\operatorname{red}}\rightarrow[X/G]$ one can pushforward the sheaves along the inclusion and one can make the following identifications
(68) $H^{1}([X/G]_{\operatorname{red}},\mathcal{F}^{i})=H^{1}([X/G],\iota_{*}\mathcal{F}^{i})=H^{1}_{G}(X,\iota_{*}\mathcal{F}^{i})=0,$
where the last identification comes from the previous theorem. All obstruction classes have to vanish, therefore $[X/G]$ is graded. ∎
As a consequence of gradedness, the superstack can also be written in the following form:
(69) $[X/G]\cong([X/G]_{\operatorname{red}},\wedge^{\bullet}\mathcal{E}),$
where $\mathcal{E}\coloneqq\mathcal{J}/\mathcal{J}^{2}$. As a consequence, we get an easier description of the Berezinian bundle
(70) $\operatorname{Ber}(\mathcal{T}_{[X/G]_{sm}})=\operatorname{det}(\mathcal{T}_{[X/G]_{sm}})\otimes\operatorname{det}^{-1}(\wedge^{top}\mathcal{E})$
Corollary 5.11.
The smooth locus character stack $\mathfrak{X}_{sm}$ is graded.
Given a point $x:\mathrm{Spec}(\mathbb{C})\rightarrow\mathfrak{X}$, we can consider the tangent complex at the point $x$, which is given by the two-term complex
(71) $[\mathfrak{g}\longrightarrow\mathcal{T}_{X,x}].$
Along the smooth locus, we obtain the tangent space
(72) $\mathcal{T}_{X,x}/\mathfrak{g}.$
This object will have the expected dimension of $(2g-2)\mathrm{dim}(\mathrm{OSp}(1|2))$ and for the details on the explicit calculations we will refer the interested reader to A.2.
Appendix A Computations
This appendix provides a detailed derivation of the dimension of the tangent space of the representation variety $\mathcal{T}_{X,x}$. We conclude that the character stack has the expected dimension for a super Riemann surface of genus $g$. The classical proofs can be adapted with minimal modification from the literature, which we will be pointing towards for each statement. The necessary modification one has to do to adapt these results to the supergeometric case is replace the usual dual numbers, by the so-called super dual numbers, which refers to the superalgebra $k[\varepsilon,\eta]/(\varepsilon^{2},\varepsilon\eta)$, where $\mathrm{deg(\varepsilon)=0}$ and $\mathrm{deg}(\eta)=1$. We begin by proving the following lemma.
Lemma A.1 ([21]).
Let $(A,\Delta)$ be a super Hopf algebra over k and let $I$ denote the kernel of the co-identity map $\epsilon$, then $A\cong k\oplus I$ as k-modules.
Proof.
The co-identity map is a morphism of super algebras, so in particular it is a morphism of super vector spaces, hence we get the following sequence
[…]
This is a sequence of super vector spacecs, so it splits and we get the wanted isomophism. ∎
The next theorem was originally proven in [17]. However, we will follow the more detailed presentation of [27].
Theorem A.2.
Let $R(\Gamma,G)$ be the coordinate ring of the representation variety $\mathrm{Hom}(\Gamma,G)$ for $\Gamma$ a finitely generated group and $G$ an algebraic supergroup. Then the Zariski tangent space at a point $\rho$ is isomorphic to $Z(\Gamma,Ad_{\rho})$.
Proof.
We can now describe a morphism from the supergroup $G(\mathbb{C}/[\varepsilon,\eta])$ to the Lie algebra $\mathfrak{g}$. Define
$\sigma(g)=g\pi(g)^{-1}$
where $\pi$ is a morphism of supergroups $G(\mathbb{C}[\varepsilon,\eta](\varepsilon^{2},\eta\varepsilon))\rightarrow G(\mathbb{C})$ induced by $\mathbb{C}[\varepsilon,\eta](\varepsilon^{2},\eta\varepsilon)\rightarrow\mathbb{C}$, where we send both variables to zero. By applying $\pi$ again, we can check if it is an element of the Lie algebra:
$\pi(\sigma(g))$ $=\pi((g\pi(g)^{-1}))$
$=\pi(g)\pi^{2}(g^{-1})$
$=e$
We can now compute the following
$\sigma(g_{1}g_{2})$ $=g_{1}g_{2}\pi(g_{1}g_{2})^{-1}$
$=g_{1}g_{2}\pi(g_{2})^{-1}\pi(g_{1})^{-1}$
$=g_{1}(\pi(g_{1})^{-1}\cdot\pi(g_{1}))\sigma(g_{2})\pi(g_{1})^{-1}$
$=\sigma(g_{1})\cdot Ad_{\pi(g_{1})}\sigma(g_{2})$
And we are allowed to switch to additive notation and we get the 1-cocycle condition for group cohomology
$\sigma(gh)=\sigma(g)+Ad_{\pi(g)}\sigma(h)$
So it is a cycle in $Z^{1}(\Gamma,\mathfrak{g})$.
We now also get a map from $\Gamma\rightarrow\mathfrak{g}$, by using the universal property
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\mathfrak{g}$$\scriptstyle{\rho_{U}}$$\scriptstyle{\psi_{f}}$$\scriptstyle{G(f)}$$\scriptstyle{\sigma}$
We will call this map $\Sigma_{f}\coloneqq\sigma\circ\psi_{f}$, which fulfills the cocycle condition, because $\psi_{f}$ is a group homomorphism and the properties of $\sigma$. Consider the diagram
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$G(\mathbb{C})$$\scriptstyle{\rho_{U}}$$\scriptstyle{\rho}$$\scriptstyle{G(f)}$$\scriptstyle{G(h_{\rho})}$
Let $f\in T_{\rho}G=\{f|R(\Gamma,G)\xrightarrow{f}k[\varepsilon,\eta]\xrightarrow{\pi}k,\pi\circ f=h_{\rho}\}$ then we get a morphism
$\Psi_{\rho}:\ T_{\rho}\mathrm{Hom}(\Gamma,G)$ $\longrightarrow Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$
$f$ $\longmapsto\Sigma_{f}$
Now let us also define an inverse. Let $\Sigma\in Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$
$\phi_{\Sigma}:\ \Gamma$ $\longrightarrow G(\mathbb{\varepsilon,\eta})$
$\gamma$ $\longmapsto\Sigma(\gamma)\cdot\rho(\gamma)$
We calculate:
$\phi_{\Sigma}(\gamma\delta)$ $=\Sigma(\gamma\delta)\cdot\rho(\gamma)$
$=\Sigma(\gamma)\cdot Ad_{\rho(\gamma)}\Sigma(\delta)\cdot\rho(\gamma\delta)$
$=\Sigma(\gamma)\rho(\gamma)\Sigma(\delta)\rho(\gamma)^{-1}\rho(\gamma)\rho(\delta)$
$=\Sigma(\gamma)\rho(\gamma)\Sigma(\delta)\rho(\delta)$
$=\phi_{\Sigma}(\gamma)\phi_{\Sigma}(\delta)$
So this is indeed a group homomorphism. By the universal property we get a morphism
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\scriptstyle{\rho_{U}}$$\scriptstyle{\phi_{\Sigma}}$$\scriptstyle{G(f_{\Sigma})}$
Recall that $\mathfrak{g}\cong ker(\pi)$ and $\rho(\gamma)\in G(\mathbb{C})$. It follows that $\pi\circ\phi_{\Sigma}=\rho$, because
$\pi(\phi_{\Sigma}(\gamma))=\pi(\Sigma(\gamma\cdot\rho(\gamma)))=\rho(\gamma).$
So, we get the map
$\Phi:\ Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})$ $\longrightarrow T_{\rho}\mathrm{Hom}(\Gamma,G)$
$\Sigma$ $\longmapsto f_{\Sigma}$
The only thing remaining is that we have to prove, that $\Psi_{\rho},\Phi_{\rho}$ are inverse to each other. Applying $\Phi_{\rho}(\Psi_{\rho}(f))=f_{\Sigma_{f}}$. By the universal property
$\Gamma$$G(R(\Gamma,G))$$G(\mathbb{C}[\varepsilon,\eta])$$\scriptstyle{\rho_{U}}$$\scriptstyle{\phi_{\Sigma_{f}}}$$\scriptstyle{G(f_{\Sigma_{f}})}$
we can now calculate
$\phi_{\Sigma_{f}}(\gamma)$ $=\Sigma_{f}(\gamma)\cdot\rho(\gamma)$
$=(\sigma\circ G(f)\circ\rho_{U})(\gamma)\cdot\rho(\gamma)$
$=(G(f)\circ\rho_{U}(\gamma))\cdot((\pi\circ G(f)\circ\rho_{U})(\gamma))^{-1}\cdot\pi\circ G(f)\circ\rho_{U})(\gamma)$
$\phi_{\Sigma_{f}}(\gamma)$ $=G(f)\circ\rho_{U}(\gamma)$
By uniqueness of the universal property, we get $\phi_{\Sigma_{f}}=\psi_{f}$. It follows that $G(f)=G(f_{\Sigma_{f}})$ so we get $f=f_{\Sigma_{f}}$. The other direction goes as follows:
$(\Psi_{\rho}\circ\Phi_{\rho})(\Sigma)$ $=\sigma\circ\psi_{f_{\Sigma}}$
$=\sigma\circ G(f_{\Sigma}\circ\rho_{U})$
$=(G(f_{\Sigma})\circ\rho_{U})(\gamma)((\pi\circ G(f_{\Sigma})\circ\rho_{U})(\gamma))^{-1}$
$=\phi_{\Sigma}(\gamma)\cdot\rho(\gamma)(\pi(\phi_{\Sigma}(\gamma)))^{-1}$
$=\Sigma(\gamma)\cdot\rho(\gamma)(\pi(\Sigma(\gamma)\cdot\rho(\gamma)))^{-1}$
$=\Sigma(\gamma)\cdot\rho(\gamma)\cdot e\cdot\rho(\gamma)^{-1}$
$=\Sigma(\gamma)$
So indeed the two morphisms are inverse to each other and we get the isomorphism
$Z^{1}(\Gamma,\mathfrak{g}_{Ad_{\rho}})\cong T_{\rho}\mathrm{Hom}(\Gamma,G)$
The next proposition, was originally proven in [13] and we will follow the explicit exposition given in [19].
Proposition A.3.
Let $X$ be a Riemann surface of genus $g$ and $G\in\mathrm{Grp(Sch)}$, such that it admits a symmetric, bilinear, non-degenerate pairing. Then the dimension of the 1-cocycles is given by
(73) $\mathrm{dim}Z^{1}(\Gamma,Ad_{\rho})=(2g-1)\mathrm{dim}(G)+\mathrm{dim}(C_{G}(\pi_{1}(X))).$
Proof.
We give a sketch of the proof by providing the key arguments, which should convince the reader. First of all, since $X$ is an Eilenberg-MacLane space $K(\pi_{1}(X),1)$ computing group cohomology is the same as computing de Rham cohomology with local coefficients, which lets us identify
(74) $H^{*}(\pi_{1}(X),Ad_{\rho})\cong H^{*}_{dR}(X,E_{\rho}),$
so the cohomology vanishes in degrees larger than $2$. Furthermore, we have the Euler characteristic formula,
(75) $\chi(X,\rho)=\mathrm{dim}(H^{0}(\pi_{1}(X),Ad_{\rho}))-\mathrm{dim}(H^{1}(\pi_{1}(X),Ad_{\rho}))+\mathrm{dim}(H^{2}(\pi_{1}(X),Ad_{\rho})).$
This is independent of $\rho$, because when defining group cohomology, the twisting by $\rho$ is only realized inside of the differentials of the chain complex, so the chain complex itself does not see the twisting. For the trivial representation $\rho$ the Euler characteristic turns into
(76) $\chi(X,\mathrm{triv})=\chi(X)\mathrm{dim}(G).$
By Poincaré duality and the non-degeneracy of the Killing form, we obtain
(77) $H^{2}(\pi_{1}(X),Ad_{\rho})\overset{PD}{\cong}H^{0}(\pi_{1}(X),Ad_{\rho}^{\vee})^{\vee}\overset{Ad_{\rho}^{\vee}\cong Ad_{\rho}}{\cong}H^{0}(\pi_{1}(X),Ad_{\rho}).$
It is true, that $H^{0}(\pi_{1}(X),Ad_{\rho})$ is the space of $Ad(\rho)$-invariant elements and also that $\mathrm{dim}(B^{1}(\pi_{1}(X),Ad_{\rho}))=\mathrm{dim}(G)-\mathrm{dim}(C(\rho))$. Combining these two facts we obtain f
(78) $\begin{split}\mathrm{dim}(Z^{1}(\pi_{1}(X),Ad_{\rho}))&=\mathrm{dim}(H^{1}(\pi_{1}(X),Ad_{\rho}))+\mathrm{dim}(B^{1}(\pi_{1}(X),Ad_{\rho}))\\ &=\chi(X)\mathrm{dim}(G)+\mathrm{dim}(G)-\mathrm{dim}(C(\rho))\\ &=(2g-2)\mathrm{dim}(G)+\mathrm{dim}(G)-\mathrm{dim}(C(\rho))\\ &=(2g-1)\mathrm{dim}(G)+\mathrm{dim}(C(\rho))\end{split}$
We note that $C(\rho)=C(G)$ minimizes the dimension. ∎
As a corollary we obtain our wanted dimension of the tangent space.
Corollary A.4.
The dimension of the tangent space of a smooth point of $\mathrm{Rep}(\pi_{1}(X),\mathrm{OSp}(1|2))$ is $(2g-1)\mathrm{dim}(\mathrm{OSp}(1|2))$.
Proof.
A point is smooth if the dimension of the scheme and the Zariski tangent space coincide. Furthermore, the center of $\mathrm{OSp}(1|2)$ is given by $\{\pm\mathrm{id}\}$, thus the dimension of the center is 0 dimensional. By (A.3) the claim follows. ∎